### v1

# Fractal transference principles for subsets of $\mathbb{N}^{d}$ of positive density

Abstract
We establish a multidimensional fractal transference principle for digit-restricted sets associated with subsets of $\mathbb{N}^{d}$, extending the one-dimensional framework of Nakajima–Takahasi [24]. We develop general Hausdorff-dimension tools via the singular value potential $\phi^{s}(\mathbf{a})$ and the multivariate Dirichlet series $\zeta_{S}(\bm{\sigma})=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\sigma_{j}}$. Let $s_{\ast}:=\inf\{s>0:\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty\}$ and $\Lambda_{S}:=\inf\{\sigma_{1}+\cdots+\sigma_{d}:\zeta_{S}(\bm{\sigma})<\infty\}$. We obtain $\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}$, where $\mathcal{E}_{S}\subset(0,1)^{d}$ denotes the set of points whose continued-fraction digit vectors lie in $S$ and whose coordinates escape (i.e. $a_{n}(x_{j})\to\infty$ for each $j$), and $s_{\ast}=\tfrac{1}{2}\Lambda_{S}$ for uniformly KK–balanced $S$. In particular, if $S\subset\mathbb{N}^{d}$ has positive upper (or upper Banach) density then $\dim_{H}(\mathcal{E}_{S})=d/2$. On the combinatorial side, the transference principle ensures that translation-invariant configurations forced at positive density, including multidimensional Szemerédi patterns, persist inside the induced fractal digit sets.
1 Introduction
In this paper we study digit-restricted continued-fraction fractals through the lens of fractal transference principles. Roughly speaking, we show that density phenomena in $\mathbb{N}^{d}$ can be encoded into Hausdorff-typical digit sets for continued fractions without loss of Hausdorff dimension. Along the way we obtain sharp Hausdorff-dimension bounds for the resulting digit-restricted limit sets.
Regular continued fractions give a canonical symbolic model for the Gauss map and are a cornerstone of metric Diophantine approximation; see, e.g., [17, 4, 15, 28]. Restricting the allowed digits in the continued-fraction expansion leads to Cantor-like limit sets in $[0,1]$, whose size is naturally measured by Hausdorff dimension. This circle of ideas originates in the work of Good and Hirst [10, 14, 13] and has been substantially developed since then (see, e.g., [6] and references therein). Our results fit into this tradition, but with an additional combinatorial input coming from density questions on digit sets.
A convenient modern approach treats digit-restricted continued-fraction sets via conformal iterated function systems and thermodynamic/symbolic methods [22, 23, 29, 27]. This framework also highlights that, beyond Hausdorff dimension, finer invariants such as Hausdorff measure at the critical dimension and packing dimension can be subtle, especially for infinite alphabets. For example, Mauldin and Urbański [22] characterized the packing dimension of limit sets for general conformal iterated function systems and applied this to continued fractions with restricted digits, relating geometric size to arithmetic density properties of the allowed entries; see also [23]. In particular, the critical Hausdorff measure may behave delicately, and packing-type substitutes and arithmetic criteria often enter (cf. [22, 23]).
A basic theme in this area is that arithmetic growth of the allowed digits often controls the fractal size of the corresponding limit set. To illustrate this in dimension one, let $B$ be an infinite increasing sequence of positive integers $b_{1}<b_{2}<\cdots$, and let $\tau$ be the exponent of convergence of the series $\sum_{n=1}^{\infty}b_{n}^{-s}$, namely
$\tau:=\inf\Bigl\{s\geq 0:\ \sum_{n=1}^{\infty}b_{n}^{-s}<\infty\Bigr\}.$
For the limit set
$\bigl\{x\in[0,1]:\ a_{n}(x)\in B\ \text{for all }n\geq 1\ \text{and}\ a_{n}(x)\to\infty\ \text{as }n\to\infty\bigr\},$
[13] gives the upper bound $\tau/2$ for its Hausdorff dimension, and [30] shows this bound is attained. Related dimension problems for limsup-type digit conditions have also been studied; see, for example, Wang–Wu [31].
Digit restrictions are also closely tied to arithmetic and applications. Bounded-digit sets connect to rational-approximation problems such as Zaremba-type questions and to transfer-operator methods; see, e.g., [12, 3, 18]. For example, Zaremba’s conjecture asks whether every positive integer can occur as the denominator of a rational number whose continued-fraction digits are uniformly bounded by an absolute constant, and Bourgain–Kontorovich [3] established this for a set of denominators of density one, highlighting links to low-discrepancy constructions in numerical integration and pseudorandom number generation. Continued-fraction algorithms also play a renormalization role beyond the Gauss map: Berthé–Steiner–Thuswaldner [2] construct symbolic codings for toral translations using multidimensional continued-fraction algorithms realized by sequences of substitutions, obtaining Sturmian-type low complexity and multiscale bounded remainder sets.
These examples emphasize that digit restrictions encode arithmetic information in a geometric way. In this paper, however, we work with the coordinatewise continued fraction expansion and impose digit constraints through subsets of $\mathbb{N}^{d}$.
On the additive-combinatorial side, Szemerédi-type density theorems show that positive-density subsets of $\mathbb{N}$ and $\mathbb{N}^{d}$ contain rich translation-invariant configurations [26, 8, 9], with polynomial extensions due to Bergelson–Leibman [1]. A direct bridge between these themes is provided by fractal transference principles: Nakajima–Takahasi [24] established a fractal transference principle in dimension one: translation-invariant density statements can be transferred into digit sets inside continued-fraction fractals without loss of Hausdorff dimension, under an injectivity constraint on digits. The present paper develops higher-dimensional analogues for digit vectors in $\mathbb{N}^{d}$, and complements them with sharp upper bounds for $\dim_{H}(\mathcal{E}_{S})$ in terms of arithmetic growth data of $S\subset\mathbb{N}^{d}$.
We now fix notation and explain how digit vectors in $\mathbb{N}^{d}$ encode subsets of $(0,1)^{d}$.
1.1 Coordinatewise continued fractions and digit sets
We recall the coordinatewise continued-fraction coding and introduce the digit sets that link subsets of $\mathbb{N}^{d}$ with digit-restricted fractals in $(0,1)^{d}$.
Every irrational $x\in(0,1)$ admits a unique infinite regular continued fraction expansion $x=[0;a_{1}(x),a_{2}(x),\dots]$, where $a_{n}(x)\in\mathbb{N}$ are the partial quotients. For $\mathbf{x}=(x_{1},\dots,x_{d})\in((0,1)\setminus\mathbb{Q})^{d}$ we use the coordinatewise expansion and define the digit vector at time $n$ by
$\mathbf{a}_{n}(\mathbf{x}):=\bigl(a_{n}(x_{1}),\dots,a_{n}(x_{d})\bigr)\in\mathbb{N}^{d}.$
While the coordinate dynamics are independent, digit restrictions will couple the coordinates through a subset $S\subset\mathbb{N}^{d}$.
In our transference arguments it is convenient to work in a regime where digits are seen only once in time, so that a point naturally generates an infinite digit set. In dimension $d=1$ we consider
$E_{1}:=\{x\in(0,1)\setminus\mathbb{Q}:\ a_{m}(x)\neq a_{n}(x)\ \text{for all }m\neq n\}.$
The study of $E_{1}$ was suggested by P. Erdős; see Ramharter [25, p. 11]. Ramharter [25] proved that $\dim_{H}(E_{1})=\tfrac{1}{2}$; see also Jordan–Rams [16]. For comparison, Good’s landmark theorem [10, Theorem 1] asserts that the escaping set
$\{x\in(0,1)\setminus\mathbb{Q}:\ a_{n}(x)\to\infty\ \text{as }n\to\infty\}$
also has Hausdorff dimension $\tfrac{1}{2}$.
For $d\geq 1$ we write
$\mathcal{E}^{\mathrm{inj}}:=E_{1}^{d}=\Bigl\{\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}:a_{m}(x_{j})\neq a_{n}(x_{j})\ \text{for all }m\neq n,\ 1\leq j\leq d\Bigr\}.$
As a corollary of Theorem 1.6, we have $\dim_{H}(\mathcal{E}^{\mathrm{inj}})=d/2$. Although $\mathcal{E}^{\mathrm{inj}}$ is a Cartesian product, the corresponding product upper bound is not automatic for general fractal sets; in our setting it follows from a classical product estimate (see Lemma 3.14 in Section 3 and Marstrand [19]).
Similar product phenomena occur in other symbolic codings. For example, Cheng–Zhang [5, Theorem 1.4] study exact-approximation sets arising from $\beta$-expansions and obtain a sharp formula for the Hausdorff dimension of their Cartesian products when $d\geq 2$. At the level of critical Hausdorff measure, even very classical self-similar products can be delicate: Guo–Jones [11] provide improved upper and lower bounds for the critical Hausdorff measure of Cartesian products of the middle-third Cantor set, extending what was previously known in dimension two.
For each $\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}$ we define the associated digit set
$D(\mathbf{x}):=\{\mathbf{a}_{n}(\mathbf{x}):n\geq 1\}\subset\mathbb{N}^{d},$
and note that $D(\mathbf{x})$ has no repetitions whenever $\mathbf{x}\in\mathcal{E}^{\mathrm{inj}}$. This provides the basic interface between combinatorics on $\mathbb{N}^{d}$ and digit-restricted fractal subsets of $(0,1)^{d}$.
On $\mathbb{N}^{d}$ we use the standard Følner boxes $Q_{N}:=[1,N]^{d}\cap\mathbb{N}^{d}$ and their translates $Q_{N}(\mathbf{v}):=\mathbf{v}+Q_{N}$. For $S\subset\mathbb{N}^{d}$ we write
$\overline{d}(S):=\limsup_{N\to\infty}\frac{|S\cap Q_{N}|}{|Q_{N}|},\qquad d^{\ast}(S):=\limsup_{N\to\infty}\sup_{\mathbf{v}\in\mathbb{N}^{d}}\frac{|S\cap Q_{N}(\mathbf{v})|}{|Q_{N}|}$
for the upper density and upper Banach density, respectively.
Finally, for an infinite $S\subset\mathbb{N}^{d}$ we write $\widetilde{\mathcal{E}}_{S}$ and $\mathcal{E}_{S}$ for the digit-restricted sets introduced in Subsection 1.4. When we need the digit vectors to be injective in time, we set
$\mathcal{E}_{S}^{\mathrm{vec}}:=\Bigl\{\mathbf{x}\in\widetilde{\mathcal{E}}_{S}:\ \mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})\ (n\neq m),\ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\Bigr\}.$
1.2 Fractal transference principles on $\mathbb{N}^{d}$
Our first main theorem is an $\mathbb{N}^{d}$-analogue of the one-dimensional transference principle of Nakajima–Takahasi [24, Theorem 1.1]. Informally, it asserts that one can choose Hausdorff-typical points in the escaping, vector-injective regime $\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ so that their digit sets reflect the density of a prescribed subset $S\subset\mathbb{N}^{d}$.
Before stating our results, we recall the one-dimensional theorem of Nakajima–Takahasi [24, Theorem 1.1]. In dimension $d=1$, Nakajima–Takahasi [24, Theorem 1.1] proved that if $S\subset\mathbb{N}$ has $\overline{d}(S)>0$ (resp. $d^{\ast}(S)>0$), then there exist sets $E_{S}$ (resp. $F_{S}$) with
$E_{S}\subset\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}\quad(\text{resp.\ }F_{S}\subset\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}),\qquad\dim_{H}E_{S}=\dim_{H}F_{S}=\dim_{H}(\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}})=\frac{1}{2},$
such that the following two statements hold.
  1. (i)
    $\overline{d}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{x\in E_{S}}\{a_{n}(x)\}\cap S\Bigr)=\overline{d}(S)\qquad(\text{resp.\ }d^{\ast}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\Bigr)=d^{\ast}(S));$
  2. (ii)
    $\dim_{H}\Bigl\{x\in\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}:\ \overline{d}(\{a_{n}(x):n\geq 1\}\cap S)=\overline{d}(S)\Bigr\}=\dim_{H}(\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}})$
    (resp. with $d^{\ast}$ in place of $\overline{d}$).
We now state an $\mathbb{N}^{d}$-analogue of this fractal transference principle.
Theorem 1.1 (Fractal transference principle on $\mathbb{N}^{d}$).
Let $S\subset\mathbb{N}^{d}$.
  • (a)
    If $\overline{d}(S)>0$, then there exists a set $E_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ such that
    $\dim_{H}E_{S}=\dim_{H}(\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}})=\frac{d}{2}$
    and
    $\overline{d}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in E_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap S\Bigr)=\overline{d}(S).$ (1)
    In particular,
    $\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}:\ \overline{d}\bigl(\{\mathbf{a}_{n}(\mathbf{x}):n\geq 1\}\cap S\bigr)=\overline{d}(S)\Bigr\}=\frac{d}{2}.$ (2)
  • (b)
    If $d^{\ast}(S)>0$, then there exists a set $F_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ such that
    $\dim_{H}F_{S}=\dim_{H}(\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}})=\frac{d}{2}$
    and
    $d^{\ast}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in F_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap S\Bigr)=d^{\ast}(S).$ (3)
    In particular,
    $\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}:\ d^{\ast}\bigl(\{\mathbf{a}_{n}(\mathbf{x}):n\geq 1\}\cap S\bigr)=d^{\ast}(S)\Bigr\}=\frac{d}{2}.$ (4)
Remark 1.2.
When $d=1$, (1)–(4) recover the fractal transference principle of Nakajima–Takahasi [24, Theorem 1.1]. Our proof follows the same surgery–and–elimination paradigm, but several points are genuinely high–dimensional: one must enforce coordinatewise divergence ($\min_{1\leq i\leq d}a_{n}(x_{i})\to\infty$) and vector–injectivity ($\mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})$ for $n\neq m$), and the dimension analysis of the model families is anisotropic, governed by dd–dimensional cylinder geometry rather than one–dimensional lengths.
Consequently, any translation–invariant configuration property that holds in every subset of $\mathbb{N}^{d}$ of positive (Banach) density transfers to $\{\mathbf{a}_{n}(\mathbf{x}):n\geq 1\}\cap S$ for $\dim_{H}$–typical $\mathbf{x}\in\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$. For instance, combining Theorem 1.1 with the polynomial Szemerédi theorem of Bergelson–Leibman [1] yields the following.
Corollary 1.3 (Polynomial configurations in digit sets).
Fix $r\in\mathbb{N}$. Let $P:\mathbb{Z}^{r}\to\mathbb{Z}^{d}$ be a polynomial mapping with $P(\mathbf{0})=\mathbf{0}$, and let $F\subset\mathbb{Z}^{r}$ be finite. Let $S\subset\mathbb{N}^{d}$.
  • (a)
    If $\overline{d}(S)>0$, then there exists $E_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ with $\dim_{H}E_{S}=d/2$ such that the set
    $\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in E_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap S$
    has positive upper density in $\mathbb{N}^{d}$. Consequently, by the polynomial Szemerédi theorem of Bergelson–Leibman [1], there exist $n\in\mathbb{N}$ and $\mathbf{u}\in\mathbb{N}^{d}$ such that for every $\mathbf{x}\in E_{S}$ one has
    $\mathbf{u}+P(nF)\subset\{\mathbf{a}_{m}(\mathbf{x}):m\geq 1\}\cap S.$
  • (b)
    If $d^{\ast}(S)>0$, then there exists $F_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ with $\dim_{H}F_{S}=d/2$ such that the set
    $\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in F_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap S$
    has positive upper Banach density in $\mathbb{N}^{d}$. Consequently, by Bergelson–Leibman [1], there exist $n\in\mathbb{N}$ and $\mathbf{u}\in\mathbb{N}^{d}$ such that for every $\mathbf{x}\in F_{S}$ one has
    $\mathbf{u}+P(nF)\subset\{\mathbf{a}_{m}(\mathbf{x}):m\geq 1\}\cap S.$
1.3 Relative fractal transference on $\mathbb{N}^{d}$
We also develop a relative version of Nakajima–Takahasi [24, Theorem 1.4], where a sparse ambient set $S\subset\mathbb{N}^{d}$ plays the role of a background set. For sharp upper bounds it is convenient to control anisotropy: for $K\geq 1$ let
$\mathcal{C}_{K}:=\Bigl\{\mathbf{a}\in\mathbb{N}^{d}:\ \frac{\max_{j}a_{j}}{\min_{j}a_{j}}\leq K\Bigr\},$
and call $S$ uniformly $K$-balanced if $S\subset\mathcal{C}_{K}$.
To quantify sparsity, we say that $S$ has polynomial density of exponent $\alpha\geq 1$ if
$|S\cap Q_{N}|\asymp\frac{N^{d/\alpha}}{(\log N)^{\beta}},$
where for two nonnegative quantities depending on $N$ we write $A_{N}\asymp B_{N}$ if there exist constants $c,C>0$ such that $c\,B_{N}\leq A_{N}\leq C\,B_{N}$ for all sufficiently large $N$.
Before stating our higher-dimensional result, we recall the one-dimensional theorem of Nakajima–Takahasi [24, Theorem 1.4]. In dimension $d=1$, they they showed that if $S\subset\mathbb{N}$ is infinite with polynomial density of exponent $\alpha\geq 1$ and $A\subset S$ satisfies $\overline{d}(A\mid S)>0$, then there exists a set $E_{S,A}\subset\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}$ with
$\{a_{n}(x):n\geq 1\}\subset S\ \text{ for all }x\in E_{S,A},\qquad\dim_{H}E_{S,A}=\dim_{H}\bigl(\mathcal{E}_{S}^{\mathrm{vec}}\bigr)=\frac{1}{2\alpha},$
such that the following two statements hold.
  1. (i)
    $\overline{d}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\ \bigm|\ S\Bigr)=\overline{d}(A\mid S);$
  2. (ii)
    $\dim_{H}\Bigl\{x\in\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}:\ \{a_{n}(x):n\geq 1\}\subset S,\ \overline{d}(\{a_{n}(x):n\geq 1\}\cap A\mid S)=\overline{d}(A\mid S)\Bigr\}=\dim_{H}\bigl(\mathcal{E}_{S}^{\mathrm{vec}}\bigr).$
We now state an $\mathbb{N}^{d}$-analogue of this relative fractal transference principle.
Theorem 1.4 (Relative fractal transference on $\mathbb{N}^{d}$).
Let $S\subset\mathcal{C}_{K}\subset\mathbb{N}^{d}$ be infinite with polynomial density of exponent $\alpha\geq 1$, and let $A\subset S$ satisfy $\overline{d}(A\mid S)>0$. Then
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=\frac{d}{2\alpha},$
and moreover
$\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{S}^{\mathrm{vec}}:\ \overline{d}\bigl(D(\mathbf{x})\cap A\mid S\bigr)=\overline{d}(A\mid S)\Bigr\}=\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}}).$
Remark 1.5.
The uniform balance hypothesis excludes extremely anisotropic digit vectors and places the polynomial-growth assumption in a regime where our covering argument applies cleanly and yields sharp upper bounds; without it, polynomial density alone need not be sufficient for our method.
  1. (i)
    When $d=1$ this hypothesis is automatic, so Theorem 1.4 recovers the one-dimensional result of Nakajima–Takahasi [24, Theorem 1.4] without any additional geometric restriction.
  2. (ii)
    In higher dimensions it is natural to ask how far the balance condition can be relaxed. For instance, in strongly anisotropic situations where admissible digits grow at very different rates across coordinates, it is unclear how the distribution of $S$ among directions should affect the dimension of $\mathcal{E}_{S}^{\mathrm{vec}}$, and whether one should expect product-type behaviour reflecting only partial independence between directions.
1.4 Hausdorff dimension of digit-restricted sets
Finally, we record our dimension estimates for digit-restricted limit sets. They give a universal upper bound and, under mild arithmetic hypotheses on $S\subset\mathbb{N}^{d}$, sharp bounds expressed in terms of a singular-value potential.
Let $S\subset\mathbb{N}^{d}$ and define
$\widetilde{\mathcal{E}}_{S}:=\{\mathbf{x}\in(0,1)^{d}:\ \mathbf{a}_{n}(\mathbf{x})\in S\ \forall n\geq 1\},\qquad\mathcal{E}_{S}:=\{\mathbf{x}\in\widetilde{\mathcal{E}}_{S}:\ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\}.$ (5)
We also recall the vector-injective subregime
$\mathcal{E}_{S}^{\mathrm{vec}}:=\Bigl\{\mathbf{x}\in\widetilde{\mathcal{E}}_{S}:\ \mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})\ (n\neq m),\ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\Bigr\}.$
Accordingly, all dimension estimates are tail statements. Since modifying finitely many initial digit vectors does not change Hausdorff dimension (Lemma 3.7), the estimates are governed by the tail geometry of cylinders.
To implement the covering argument in Section 3, we encode the anisotropic cylinder geometry by a singular-value potential. For $\mathbf{a}=(a_{1},\dots,a_{d})\in\mathbb{N}^{d}$ let $a_{(1)}\leq\cdots\leq a_{(d)}$ denote the nondecreasing rearrangement. For $s\in(0,d]$ and $k<s\leq k+1$ we set
$\phi^{s}(\mathbf{a}):=\Bigl(\prod_{j=1}^{k}a_{(j)}^{-2}\Bigr)\,a_{(k+1)}^{-2(s-k)}.$
This potential yields a cylinder covering criterion and the associated critical threshold
$s_{\ast}:=\inf\Bigl\{s>0:\ \sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty\Bigr\},$
in particular, $s_{\ast}$ is the critical exponent governing our upper bounds. We refer to Section 3 for the covering argument and its equivalent Dirichlet-series formulation.
Theorem 1.6 (Universal upper bound and sharpness).
For every infinite $S\subset\mathbb{N}^{d}$, one has $\dim_{H}(\mathcal{E}_{S})\leq d/2$. If moreover $\overline{d}(S)>0$, then $\dim_{H}(\mathcal{E}_{S})=d/2$.
Corollary 1.7 (Polynomial growth in the balanced regime).
Assume $S\subset\mathcal{C}_{K}$ for some $K\geq 1$ and $(S\cap Q_{N})\ll N^{d/\alpha}$ for some $\alpha\geq 1$. Then $\dim_{H}(\mathcal{E}_{S})\leq d/(2\alpha)$.
Corollary 1.8 (Vector-injective subregime).
Let $S\subset\mathbb{N}^{d}$ be infinite and recall
$\mathcal{E}_{S}^{\mathrm{vec}}:=\{\mathbf{x}\in\mathcal{E}_{S}:\ \mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})\ \text{for all }n\neq m\}.$
  • (a)
    If $\overline{d}(S)>0$, then $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=d/2$.
  • (b)
    If $S\subset\mathcal{C}_{K}$ is uniformly KK–balanced and has polynomial density of exponent $\alpha\geq 1$, then $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=\dim_{H}(\mathcal{E}_{S})=d/(2\alpha)$.
The first statement of Theorem 1.6 is a universal upper bound, and the second shows sharpness in the positive-density case.
We note that the results proved in Section 3 already imply the statements recorded here, and in fact yield more general conclusions. In particular, Theorem 1.6 is the bound-and-sharpness case of Theorem 3.1, while Corollary 1.7 is a restatement of Corollary 3.5. In the uniformly KK–balanced setting, polynomial growth of $S$ translates into a corresponding drop in dimension. Finally, Corollary 1.8 combines these statements with the observation that our lower-bound constructions may be arranged within the vector-injective subregime $\mathcal{E}_{S}^{\mathrm{vec}}$.
1.5 Organization of the paper
Section 2 collects the basic tools used throughout the paper. In Subsection 2.1 we recall the required continued–fraction estimates and the geometry of coordinatewise cylinders in $(0,1)^{d}$. Subsection 2.2 fixes notation for upper density and upper Banach density on $\mathbb{N}^{d}$, together with the relative versions used later. Subsection 2.3 records the almost–Lipschitz / Hölder stability statements for elimination maps, which are the quantitative input for the surgery step.
Section 3 develops the covering framework for digit–restricted sets in $\mathbb{N}^{d}$. We formulate the singular–value potential and the associated Dirichlet–series criteria, and prove the universal upper bounds and their sharpness statements, culminating in Theorems 3.1 and 3.4 (together with the polynomial–growth consequences).
Section 4 develops higher–dimensional Moran constructions for coordinatewise continued–fraction cylinders, providing the model families and lower–bound mechanisms used later, compatible with coordinatewise divergence and, when needed, vector–injectivity.
Section 5 implements the three–stage construction on $\mathbb{N}^{d}$. Subsection 5.1 produces a relatively thin subset with good bookkeeping properties; Subsection 5.2 constructs extreme seed sets in the complement and computes their dimension; Subsection 5.3 performs the digit insertion and records the elimination estimates that transfer density information into digit sets without loss of Hausdorff dimension.
Finally, Section 6 assembles these ingredients to prove the main dimension theorems and the transference principles stated in the introduction, including the positive–density and positive–Banach–density regimes.
2 Preliminaries
This section gathers several technical lemmas that will be used repeatedly in the subsequent proofs. Their role is to provide quantitative control over density notions on $\mathbb{N}^{d}$, as well as the basic geometric estimates for coordinatewise continued–fraction cylinders.
Compared with the one–dimensional setting, additional care is required in higher dimensions to handle translations of density boxes, anisotropic growth of digit vectors, and the fact that the geometry of multidimensional cylinders is governed by the largest coordinate scale. The results collected here allow us to reduce to well–positioned and uniformly balanced regimes, and to carry out the later constructions without losing control of density or Hausdorff dimension.
2.1 Some facts on continued fractions
We briefly recall a few standard facts about regular continued fractions that will be used throughout the paper. For each $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, define the convergents $p_{n}/q_{n}$ recursively by:
$p_{-1}$ $=1,\quad p_{0}=0,\quad p_{i}=a_{i}p_{i-1}+p_{i-2}\quad\text{for }i=1,\ldots,n,$
$q_{-1}$ $=0,\quad q_{0}=1,\quad q_{i}=a_{i}q_{i-1}+q_{i-2}\quad\text{for }i=1,\ldots,n.$
The $n$-th fundamental interval is defined by:
$I(a_{1},\ldots,a_{n})=\begin{cases}\left[\dfrac{p_{n}}{q_{n}},\dfrac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}\right)&\text{if }n\text{ is even},\\[10.0pt] \left(\dfrac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}},\dfrac{p_{n}}{q_{n}}\right]&\text{if }n\text{ is odd}.\end{cases}$
This interval consists of all $x\in(0,1)$ whose regular continued fraction expansion begins with $a_{1},\ldots,a_{n}$.
The following classical estimates provide control over the geometry of these intervals. Both estimates follow from standard arguments in continued fraction theory.
Lemma 2.1.
For every $n\in\mathbb{N}$ and $(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$, the following hold.
  1. (i)
    $\frac{1}{2}\prod_{i=1}^{n}\frac{1}{(a_{i}+1)^{2}}\;\leq\;|I(a_{1},\dots,a_{n})|\;\leq\;\prod_{i=1}^{n}\frac{1}{a_{i}^{2}}.$
  2. (ii)
    For any two irrational numbers $x_{1},x_{2}\in(0,1)$ with the same first $n$ partial quotients $a_{1},\dots,a_{n}$ but $a_{n+1}(x_{1})\neq a_{n+1}(x_{2})$, there exists an absolute constant $c>0$ such that
    $|x_{1}-x_{2}|\;\geq\;\frac{c}{q_{n}^{2}}\,\frac{|a_{n+1}(x_{1})-a_{n+1}(x_{2})|}{a_{n+1}(x_{1})a_{n+1}(x_{2})}.$
Proof.
Both statements are standard consequences of the elementary geometry of continued fraction cylinders and can be verified by a direct computation. See, for instance, [10, 13, 24]. ∎
Lemma 2.2.
For every $n\geq 1$ and every word $(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$ with convergents $p_{n}/q_{n}$,
$|I(a_{1},\dots,a_{n})|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
In particular, $|I(a_{1},\dots,a_{n})|\asymp q_{n}^{-2}$ with absolute constants.
Proof.
See, for example, [10, 13]. ∎
We now extend these concepts to higher dimensions. For $\mathbf{x}=(x_{1},\ldots,x_{d})\in(0,1)^{d}$, define the digit vector at time $i$ as:
$\mathbf{a}_{i}(\mathbf{x})=(a_{i}(x_{1}),\ldots,a_{i}(x_{d}))\in\mathbb{N}^{d},$
where $a_{i}(x_{j})$ denotes the $i$-th partial quotient in the continued fraction expansion of $x_{j}$.
Definition 2.3.
For $\omega=(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})\in(\mathbb{N}^{d})^{n}$, the cylinder set $I_{\omega}$ is defined as:
$I_{\omega}=I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})=\{\mathbf{x}\in(0,1)^{d}:\mathbf{a}_{i}(\mathbf{x})=\mathbf{a}_{i}\text{ for }i=1,\ldots,n\}.$
The geometric structure of these cylinder sets is characterized by their product decomposition:
Lemma 2.4.
For $\omega=(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})\in(\mathbb{N}^{d})^{n}$, the cylinder set decomposes as:
$I_{\omega}=I^{(1)}_{\omega}\times I^{(2)}_{\omega}\times\cdots\times I^{(d)}_{\omega}$,
where each $I^{(j)}_{\omega}$ is the one-dimensional cylinder determined by $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$, that is $I^{(j)}_{\omega}=I(a_{1}^{(j)},\ldots,a_{n}^{(j)})$.
Proof.
The result follows immediately from the coordinate-wise definition:
$\mathbf{x}\in I_{\omega}\iff a_{i}(x_{j})=a_{i}^{(j)}\text{ for all }i=1,\ldots,n\text{ and }j=1,\ldots,d.$
Thus, membership in $I_{\omega}$ is equivalent to $x_{j}\in I^{(j)}_{\omega}$ for each coordinate $j$. ∎
This product structure allows us to establish diameter estimates for multidimensional cylinders:
Proposition 2.5.
Let $\omega=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})\in(\mathbb{N}^{d})^{n}$. Then there exist constants $c_{1},c_{2}>0$, depending only on $d$, such that
$c_{1}\max_{1\leq j\leq d}\ \prod_{i=1}^{n}(a_{i}^{(j)}+1)^{-2}\;\leq\;\operatorname{diam}(I_{\omega})\;\leq\;c_{2}\max_{1\leq j\leq d}\ \prod_{i=1}^{n}(a_{i}^{(j)})^{-2}.$
Proof.
By Lemma 2.4, $I_{\omega}=\prod_{j=1}^{d}I^{(j)}_{\omega}$ with $I^{(j)}_{\omega}=I(a_{1}^{(j)},\dots,a_{n}^{(j)})$. Hence
[…]
Applying Lemma 2.1(i) coordinatewise gives
$\frac{1}{2}\prod_{i=1}^{n}(a_{i}^{(j)}+1)^{-2}\;\leq\;|I^{(j)}_{\omega}|\;\leq\;\prod_{i=1}^{n}(a_{i}^{(j)})^{-2}.$
Taking the maximum over $j$ and absorbing the factors $\frac{1}{2}$ and $\sqrt{d}$ into $c_{1},c_{2}$ yields the claim. ∎
Lemma 2.6.
Let $\mathbf{x}_{1},\mathbf{x}_{2}\in(0,1)^{d}$ such that $\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})$ for $1\leq i\leq n$, but $\mathbf{a}_{n+1}(\mathbf{x}_{1})\neq\mathbf{a}_{n+1}(\mathbf{x}_{2})$. Then there exists a constant $C>0$ depending only on $d$ such that
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C\cdot\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|}{a_{n+1}^{(j)}(\mathbf{x}_{1})a_{n+1}^{(j)}(\mathbf{x}_{2})}\right),$
where $\mathcal{D}=\{1\leq j\leq d:a_{n+1}^{(j)}(\mathbf{x}_{1})\neq a_{n+1}^{(j)}(\mathbf{x}_{2})\}$.
Proof.
By the product structure from Lemma 2.4, the cylinder set $I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})$ decomposes as a Cartesian product of one-dimensional cylinder sets:
$I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})=I^{(1)}_{\omega}\times I^{(2)}_{\omega}\times\cdots\times I^{(d)}_{\omega}$,
where each $I^{(j)}_{\omega}$ is the one-dimensional cylinder determined by $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$.
For each coordinate $j\in\mathcal{D}$, the one-dimensional continued fractions $x_{1,j}$ and $x_{2,j}$ have the same first $n$ partial quotients $a_{1}^{(j)},\ldots,a_{n}^{(j)}$ but differ at the $(n+1)$-th partial quotient. By Lemma 2.1(ii), there exists an absolute constant $c_{0}>0$ such that for each $j\in\mathcal{D}$,
$|x_{1,j}-x_{2,j}|\;\geq\;\frac{c_{0}}{q_{n,j}^{2}}\,\frac{\bigl|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})\bigr|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})},$
where $q_{n,j}$ is the denominator of the $n$-th convergent associated with $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$. Since $|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\leq q_{n,j}^{-2}$ by Lemma 2.2, we may rewrite the lower bound as
$|x_{1,j}-x_{2,j}|\;\geq\;c\,|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{\bigl|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})\bigr|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})},$
for some absolute constant $c>0$ (e.g. $c=c_{0}$).
Since $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}=\max_{1\leq j\leq d}|x_{1,j}-x_{2,j}|$, we have:
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\max_{j\in\mathcal{D}}|x_{1,j}-x_{2,j}|\geq c\cdot\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|}{a_{n+1}^{(j)}(\mathbf{x}_{1})a_{n+1}^{(j)}(\mathbf{x}_{2})}\right).$
Taking $C=c$ completes the proof. Note that the constant $c$ here is the same absolute constant from the one-dimensional case and does not depend on $d$. ∎
The separation properties of these cylinder sets are inherited from the corresponding one-dimensional continued fraction geometry.
Lemma 2.7 (Sibling cylinder separation).
Let $n\geq 1$ and let $\omega,\omega^{\prime}\in(\mathbb{N}^{d})^{n}$ be distinct words with the same prefix of length n1n-1, say
$\omega=(\mathbf{a}_{1},\dots,\mathbf{a}_{n-1},\mathbf{v}),\qquad\omega^{\prime}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n-1},\mathbf{w}).$
Assume that $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$. Then $\overline{I_{\omega}}$ and $\overline{I_{\omega^{\prime}}}$ are disjoint and there exists an absolute constant $c>0$ (say $c=\tfrac{1}{4}$) such that
[…]
Proof.
By Lemma 2.4,
$I_{\omega}=\prod_{j=1}^{d}I^{(j)}_{\omega},\qquad I_{\omega^{\prime}}=\prod_{j=1}^{d}I^{(j)}_{\omega^{\prime}}$.
Choose $j_{0}$ such that $|v^{(j_{0})}-w^{(j_{0})}|\geq 2$. Let
$\alpha:=(a_{1}^{(j_{0})},\dots,a_{n-1}^{(j_{0})}),\qquad m:=v^{(j_{0})},\quad m^{\prime}:=w^{(j_{0})}.$
Then
$I^{(j_{0})}_{\omega}=I(\alpha,m),\qquad I^{(j_{0})}_{\omega^{\prime}}=I(\alpha,m^{\prime}).$
Assume w.l.o.g. $m<m^{\prime}$, so $m^{\prime}\geq m+2$. The one-dimensional depth-$n$ cylinders $\{I(\alpha,t)\}_{t\geq 1}$ form a disjoint partition of the parent interval $I(\alpha)$ (up to endpoints) in their natural order, hence the intermediate cylinder $I(\alpha,m+1)$ lies between $I(\alpha,m)$ and $I(\alpha,m^{\prime})$. Therefore
$\operatorname{dist}\bigl(I(\alpha,m),I(\alpha,m^{\prime})\bigr)\ \geq\ |I(\alpha,m+1)|.$
Write $p_{k}/q_{k}$ for the convergents associated with the prefix $\alpha$. For the word $(\alpha,t)$ one has $q_{n}(t)=tq_{n-1}+q_{n-2}$, and the endpoint formula for continued-fraction cylinders gives
$|I(\alpha,t)|=\frac{1}{q_{n}(t)\bigl(q_{n}(t)+q_{n-1}\bigr)}=\frac{1}{(tq_{n-1}+q_{n-2})\bigl((t+1)q_{n-1}+q_{n-2}\bigr)}.$
In particular, $t\mapsto|I(\alpha,t)|$ is decreasing, so
$|I(\alpha,m+1)|\ \geq\ |I(\alpha,m^{\prime})|=\min\bigl(|I(\alpha,m)|,\ |I(\alpha,m^{\prime})|\bigr).$
Hence
[…]
Finally, for Cartesian products we have
$\operatorname{dist}(I_{\omega},I_{\omega^{\prime}})\ \geq\ \operatorname{dist}\bigl(I^{(j_{0})}_{\omega},I^{(j_{0})}_{\omega^{\prime}}\bigr)$,
which yields the stated bound (with $c=1$). ∎
These geometric separation properties will be used repeatedly in the construction and analysis of higher-dimensional fractal transference arguments.
In the later construction of the seed (root) set, the $n$-th level pieces are chosen inside dd–dimensional continued–fraction cylinders. To verify the vanishing mesh condition in Definition 4.1, it is convenient to have a uniform estimate showing that one–dimensional cylinders shrink exponentially in the word length, with constants independent of the digits. A bound of the form $|I(a_{1},\dots,a_{n})|\ll\prod_{k=1}^{n}a_{k}^{-2}$ is not uniform when some digits are small, and therefore it is not suitable for deducing $\max_{Q\in\mathcal{Q}_{n}}\operatorname{diam}(Q)\to 0$ without additional growth assumptions. The following lemma is a basic property of regular continued fractions; we include a complete proof for convenience.
Lemma 2.8 (Uniform exponential decay of continued-fraction cylinders).
There exist constants $C>0$ and $\vartheta\in(0,1)$ such that for every $n\geq 1$ and every word $(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$,
$|I(a_{1},\dots,a_{n})|\leq C\,\vartheta^{n}.$
In particular one may take $\vartheta=\varphi^{-2}$, where $\varphi=(1+\sqrt{5})/2$.
Proof.
Let $p_{n}/q_{n}$ be the $n$-th convergent associated with $(a_{1},\dots,a_{n})$, with $q_{-1}=0$, $q_{0}=1$, and
$q_{n}=a_{n}q_{n-1}+q_{n-2}\qquad(n\geq 1).$
It is standard that the cylinder $I(a_{1},\dots,a_{n})$ is the open interval whose endpoints are the two rationals
$\frac{p_{n}}{q_{n}}\quad\text{and}\quad\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}$
(up to order, depending on the parity of $n$). Using the determinant identity $p_{n}q_{n-1}-p_{n-1}q_{n}=(-1)^{n-1}$, we obtain the exact length formula
$|I(a_{1},\dots,a_{n})|=\left|\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}-\frac{p_{n}}{q_{n}}\right|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
Since $a_{n}\geq 1$ for all $n$, the recursion yields $q_{n}\geq q_{n-1}+q_{n-2}$, and hence $q_{n}\geq F_{n+1}$ for all $n\geq 0$, where $(F_{m})$ is the Fibonacci sequence. By Binet’s formula,
$F_{m}=\frac{\varphi^{m}-\psi^{m}}{\sqrt{5}},\qquad\psi=\frac{1-\sqrt{5}}{2},$
so $F_{m}\geq c\,\varphi^{m}$ for all $m\geq 1$ with some absolute constant $c>0$. Combining these estimates gives
$|I(a_{1},\dots,a_{n})|\leq\frac{1}{F_{n+1}^{2}}\leq C\,\varphi^{-2n},$
which proves the claim (taking $\vartheta=\varphi^{-2}$ and adjusting $C$ if needed). ∎
Remark 2.9.
Applying Lemma 2.8 coordinatewise, if $I_{\omega}=\prod_{j=1}^{d}I^{(j)}_{\omega}$ is a dd–dimensional cylinder of word length $n$, then
$\operatorname{diam}(I_{\omega})\leq\sqrt{d}\,\max_{1\leq j\leq d}|I^{(j)}_{\omega}|\leq\sqrt{d}\,C\,\vartheta^{n}.$
In particular, any choice of cubes $Q_{\omega}\subset I_{\omega}$ at level $n$ satisfies $\max_{\omega}\operatorname{diam}(Q_{\omega})\to 0$ as $n\to\infty$.
2.2 Scaling and reduction lemmas in $\mathbb{N}^{d}$
The following lemma is used in the root set construction to control how polynomial density behaves under changes of scale.
Lemma 2.10.
Let $S\subset\mathbb{N}^{d}$ have polynomial density with exponent $\alpha\geq 1$. Then for any $t,L>1$ and all sufficiently large $n$, we have
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\asymp L^{-d/\alpha}.$
More precisely, for any $t,L>1$, there exist positive constants $C_{1}^{\prime}$ and $C_{2}^{\prime}$, independent of $L$ and $t$, such that for all sufficiently large $n$:
$C_{1}^{\prime}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq C_{2}^{\prime}L^{-d/\alpha}.$
Proof.
Since $S$ has polynomial density with exponent $\alpha$, there exist constants $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $R$:
$C_{1}\frac{R^{d/\alpha}}{(\log R)^{\beta}}\leq|S\cap Q_{R}|\leq C_{2}\frac{R^{d/\alpha}}{(\log R)^{\beta}}.$
For the upper bound:
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq\frac{C_{2}\frac{(t^{n})^{d/\alpha}}{(\log t^{n})^{\beta}}}{C_{1}\frac{(Lt^{n})^{d/\alpha}}{(\log(Lt^{n}))^{\beta}}}=\frac{C_{2}}{C_{1}}L^{-d/\alpha}\cdot\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}.$
For the lower bound:
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\geq\frac{C_{1}\frac{(t^{n})^{d/\alpha}}{(\log t^{n})^{\beta}}}{C_{2}\frac{(Lt^{n})^{d/\alpha}}{(\log(Lt^{n}))^{\beta}}}=\frac{C_{1}}{C_{2}}L^{-d/\alpha}\cdot\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}.$
Now observe that:
$\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}=\left(1+\frac{\log L}{n\log t}\right)^{\beta}\to 1\quad\text{as }n\to\infty,$
and similarly:
$\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}=\left(1-\frac{\log L}{n\log t+\log L}\right)^{\beta}\to 1\quad\text{as }n\to\infty.$
Therefore, for sufficiently large $n$, we have:
$\frac{1}{2}\leq\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}\leq 2\quad\text{and}\quad\frac{1}{2}\leq\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}\leq 2.$
Combining these estimates, we obtain:
$\frac{C_{1}}{2C_{2}}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq\frac{2C_{2}}{C_{1}}L^{-d/\alpha},$
which proves the desired asymptotic equivalence $\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\asymp L^{-d/\alpha}$. ∎
The following lemma is used in the proof of Theorem 1.1, part (b). It produces a subset $S^{\prime}\subset S$ with $d^{\ast}(S^{\prime})=d^{\ast}(S)$ and $\overline{d}(S^{\prime})=0$, witnessed along translates $Q_{N_{k}}(\mathbf{v}_{k})$ with $\min_{i}(v_{k})_{i}\to\infty$.
Lemma 2.11 (Density reduction with diverging translations).
Let $S\subset\mathbb{N}^{d}$ and set $\alpha:=d^{\ast}(S)$. Then there exist a strictly increasing sequence $(N_{k})_{k\geq 1}$ and vectors $(\mathbf{v}_{k})_{k\geq 1}\subset\mathbb{N}^{d}$, and a subset $S^{\prime}\subset S$ such that
$d^{\ast}(S^{\prime})=d^{\ast}(S)=\alpha\qquad\text{and}\qquad\overline{d}(S^{\prime})=0,$
and such that
$\frac{|S^{\prime}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}\longrightarrow\alpha\qquad(k\to\infty),$
and moreover
$\min_{1\leq i\leq d}(v_{k})_{i}\longrightarrow\infty.$
Proof.
We proceed in three steps. First, by partitioning a large cube where $S$ has density close to $\alpha$ and averaging over subcubes, we obtain cubes $Q_{N_{k}}(\mathbf{v}_{k})$ with $|S\cap Q_{N_{k}}(\mathbf{v}_{k})|/|Q_{N_{k}}|\to\alpha$ and $\min_{i}(v_{k})_{i}\to\infty$. Second, we construct $S^{\prime}\subset S$ by restricting $S$ to a rapidly separated sequence of such cubes. Third, we verify that $d^{\ast}(S^{\prime})=\alpha$, $\overline{d}(S^{\prime})=0$, and that the chosen cubes witness the required convergence.
Step 1: Far-out witnesses.
Fix a sequence $\varepsilon_{k}\downarrow 0$, for instance $\varepsilon_{k}:=1/k$. By the definition of $d^{\ast}(S)=\alpha$, for each $k$ we may choose $R_{k}\in\mathbb{N}$ and $\mathbf{u}_{k}\in\mathbb{N}^{d}$ such that
$\frac{|S\cap Q_{R_{k}}(\mathbf{u}_{k})|}{|Q_{R_{k}}|}>\alpha-\varepsilon_{k}.$
Passing to a subsequence, we may assume $R_{k}\to\infty$ and $R_{k}\geq k^{4}$.
Set $N_{k}:=k$, let $q_{k}:=\lfloor R_{k}/N_{k}\rfloor$, and define $R_{k}^{\prime}:=q_{k}N_{k}$. Then $R_{k}^{\prime}\leq R_{k}$ and $0\leq R_{k}-R_{k}^{\prime}<N_{k}$, hence
$|Q_{R_{k}}|-|Q_{R_{k}^{\prime}}|=R_{k}^{d}-(R_{k}^{\prime})^{d}\leq d(R_{k}-R_{k}^{\prime})R_{k}^{d-1}\leq dN_{k}R_{k}^{d-1}.$
Therefore,
$\frac{|S\cap Q_{R_{k}^{\prime}}(\mathbf{u}_{k})|}{|Q_{R_{k}^{\prime}}|}\geq\frac{|S\cap Q_{R_{k}}(\mathbf{u}_{k})|-(|Q_{R_{k}}|-|Q_{R_{k}^{\prime}}|)}{|Q_{R_{k}^{\prime}}|}>(\alpha-\varepsilon_{k})-\frac{dN_{k}R_{k}^{d-1}}{(R_{k}^{\prime})^{d}}.$
Since $R_{k}^{\prime}\geq R_{k}-N_{k}$ and $R_{k}\gg N_{k}$, the error term tends to 0. After passing to a further subsequence (and relabelling), we may assume that
$\frac{|S\cap Q_{R_{k}^{\prime}}(\mathbf{u}_{k})|}{|Q_{R_{k}^{\prime}}|}>\alpha-2\varepsilon_{k}\qquad\text{for all }k.$
Partition $Q_{R_{k}^{\prime}}(\mathbf{u}_{k})$ into the $q_{k}^{d}$ disjoint cubes
$Q_{N_{k}}\bigl(\mathbf{u}_{k}+N_{k}\mathbf{m}\bigr),\qquad\mathbf{m}\in\{0,1,\dots,q_{k}-1\}^{d}.$
Let $M_{k}:=k^{2}$. Call a cube $M_{k}$–interior if its translation vector $\mathbf{u}_{k}+N_{k}\mathbf{m}$ satisfies $\min_{i}\bigl((u_{k})_{i}+N_{k}m_{i}\bigr)\geq M_{k}$. Let $\mathcal{I}_{k}$ be the set of interior indices and $\mathcal{B}_{k}$ its complement. For each coordinate $i$, the number of boundary layers in the $i$th direction is at most $\lceil M_{k}/N_{k}\rceil\leq k+1$, hence
$|\mathcal{B}_{k}|\leq d(k+1)\,q_{k}^{d-1}.$
Since $R_{k}\geq k^{4}$ and $N_{k}=k$, we have $q_{k}\geq R_{k}/N_{k}-1\geq k^{3}-1$, so for all large $k$,
$\frac{|\mathcal{B}_{k}|}{q_{k}^{d}}\leq\frac{2d}{k^{2}}.$
For $\mathbf{m}\in\{0,\dots,q_{k}-1\}^{d}$ set
$\rho_{\mathbf{m}}:=\frac{|S\cap Q_{N_{k}}(\mathbf{u}_{k}+N_{k}\mathbf{m})|}{|Q_{N_{k}}|}\in[0,1].$
Averaging over all $q_{k}^{d}$ cubes gives
$\frac{1}{q_{k}^{d}}\sum_{\mathbf{m}}\rho_{\mathbf{m}}=\frac{|S\cap Q_{R_{k}^{\prime}}(\mathbf{u}_{k})|}{|Q_{R_{k}^{\prime}}|}>\alpha-2\varepsilon_{k}.$
Since $\rho_{\mathbf{m}}\leq 1$,
$\sum_{\mathbf{m}\in\mathcal{I}_{k}}\rho_{\mathbf{m}}\geq q_{k}^{d}(\alpha-2\varepsilon_{k})-|\mathcal{B}_{k}|.$
Dividing by $|\mathcal{I}_{k}|=q_{k}^{d}-|\mathcal{B}_{k}|$ and using $\frac{|\mathcal{B}_{k}|}{q_{k}^{d}}\leq\frac{2d}{k^{2}}$ yields, for all large $k$,
$\frac{1}{|\mathcal{I}_{k}|}\sum_{\mathbf{m}\in\mathcal{I}_{k}}\rho_{\mathbf{m}}\geq\alpha-2\varepsilon_{k}-\frac{2d}{k^{2}}.$
Hence there exists $\mathbf{m}_{k}\in\mathcal{I}_{k}$ such that
$\rho_{\mathbf{m}_{k}}>\alpha-2\varepsilon_{k}-\frac{2d}{k^{2}}.$
Define
$\mathbf{v}_{k}:=\mathbf{u}_{k}+N_{k}\mathbf{m}_{k}.$
Then $\min_{i}(v_{k})_{i}\geq M_{k}=k^{2}$, so $\min_{i}(v_{k})_{i}\to\infty$, and
$\frac{|S\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}=\rho_{\mathbf{m}_{k}}\longrightarrow\alpha.$
Step 2: Constructing $S^{\prime}$.
We now construct $S^{\prime}\subset S$ supported on a rapidly separated family of cubes. Fix a sequence $\varepsilon_{j}\downarrow 0$ (relabel if necessary). We construct pairwise disjoint cubes $B_{j}=Q_{T_{j}}(\mathbf{w}_{j})$ inductively and set
$S^{\prime}:=\bigcup_{j\geq 1}\bigl(S\cap B_{j}\bigr)\subset S.$
Assume $B_{1},\dots,B_{j-1}$ have been chosen and put
$H_{j}:=\max_{1\leq i\leq j-1}\bigl(\|\mathbf{w}_{i}\|_{\infty}+T_{i}\bigr)\qquad(\text{with }H_{1}:=1).$
Choose $k_{j}$ so large that
$N_{k_{j}}\geq 2H_{j},\qquad\min_{1\leq i\leq d}(v_{k_{j}})_{i}\geq 10^{j}H_{j},\qquad\frac{|S\cap Q_{N_{k_{j}}}(\mathbf{v}_{k_{j}})|}{|Q_{N_{k_{j}}}|}>\alpha-\varepsilon_{j},$
and additionally
$\frac{2^{d}d\,H_{j}}{N_{k_{j}}}\leq\frac{\varepsilon_{j}}{100}.$
Write $N_{k_{j}}=q_{j}H_{j}+r_{j}$ with $0\leq r_{j}<H_{j}$ and set $N^{\prime}_{j}:=N_{k_{j}}-r_{j}$. Then $N^{\prime}_{j}$ is a multiple of $H_{j}$ and
$|Q_{N_{k_{j}}}|-|Q_{N^{\prime}_{j}}|=N_{k_{j}}^{d}-(N^{\prime}_{j})^{d}\leq d\,r_{j}\,N_{k_{j}}^{d-1}\leq d\,H_{j}\,N_{k_{j}}^{d-1}.$
Hence
$|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|\geq|S\cap Q_{N_{k_{j}}}(\mathbf{v}_{k_{j}})|-\bigl(|Q_{N_{k_{j}}}|-|Q_{N^{\prime}_{j}}|\bigr)>(\alpha-\varepsilon_{j})N_{k_{j}}^{d}-dH_{j}N_{k_{j}}^{d-1}.$
Dividing by $|Q_{N^{\prime}_{j}}|=(N^{\prime}_{j})^{d}$ and using $(N^{\prime}_{j})^{d}\geq(N_{k_{j}}-H_{j})^{d}\geq(N_{k_{j}}/2)^{d}$ (since $N_{k_{j}}\geq 2H_{j})$, we obtain
$\frac{|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|}{|Q_{N^{\prime}_{j}}|}>(\alpha-\varepsilon_{j})-\frac{2^{d}d\,H_{j}}{N_{k_{j}}}\geq\alpha-\varepsilon_{j}-\frac{\varepsilon_{j}}{100}.$
Now partition $Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$ into disjoint cubes of side length $H_{j}$. By averaging, there exists a cube $Q_{H_{j}}(\mathbf{w}_{j})\subset Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$ such that
$\frac{|S\cap Q_{H_{j}}(\mathbf{w}_{j})|}{|Q_{H_{j}}|}\geq\frac{|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|}{|Q_{N^{\prime}_{j}}|}>\alpha-\varepsilon_{j}-\frac{\varepsilon_{j}}{100}.$
Set $T_{j}:=H_{j}$ and $B_{j}:=Q_{T_{j}}(\mathbf{w}_{j})$.
Moreover, since $B_{j}\subset Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$, we have $\min_{i}(w_{j})_{i}\geq\min_{i}(v_{k_{j}})_{i}\geq 10^{j}T_{j}$. On the other hand, the union $\bigcup_{i=1}^{j-1}B_{i}$ is contained in $Q_{H_{j}}=Q_{T_{j}}$ by the definition of $H_{j}$. Therefore $B_{j}\cap\bigcup_{i=1}^{j-1}B_{i}=\emptyset$, so the cubes $(B_{j})$ are pairwise disjoint.
Step 3: Verifying the conclusions.
For each $j$ we have $S^{\prime}\cap B_{j}=S\cap B_{j}$, hence
$d^{\ast}(S^{\prime})\geq\limsup_{j\to\infty}\frac{|S^{\prime}\cap B_{j}|}{|B_{j}|}=\limsup_{j\to\infty}\frac{|S\cap B_{j}|}{|B_{j}|}\geq\alpha.$
Since $S^{\prime}\subset S$, also $d^{\ast}(S^{\prime})\leq d^{\ast}(S)=\alpha$, hence $d^{\ast}(S^{\prime})=\alpha$.
To prove $\overline{d}(S^{\prime})=0$, let $N$ be large and let $j(N)$ be the largest index such that $B_{j(N)}\cap Q_{N}\neq\emptyset$ (if no such index exists, then $S^{\prime}\cap Q_{N}=\emptyset$). Then $B_{m}\cap Q_{N}=\emptyset$ for all $m>j(N)$, and therefore
$|S^{\prime}\cap Q_{N}|\leq\sum_{m=1}^{j(N)}|B_{m}|\leq j(N)\,T_{j(N)}^{d},$
since $T_{m}\leq T_{j(N)}$ for $m\leq j(N)$ by construction. Moreover, $B_{j(N)}\cap Q_{N}\neq\emptyset$ implies $N\geq\min_{i}(w_{j(N)})_{i}$, while $\min_{i}(w_{j(N)})_{i}\geq 10^{j(N)}T_{j(N)}$. Hence
$\frac{|S^{\prime}\cap Q_{N}|}{|Q_{N}|}\leq\frac{j(N)\,T_{j(N)}^{d}}{N^{d}}\leq\frac{j(N)\,T_{j(N)}^{d}}{\min_{i}(w_{j(N)})_{i}^{d}}\leq j(N)\,10^{-d\,j(N)}\xrightarrow[N\to\infty]{}0,$
since $j(N)\to\infty$ as $N\to\infty$. Thus $\overline{d}(S^{\prime})=0$.
Finally, define the sequences $(N_{k})$ and $(\mathbf{v}_{k})$ by
$(N_{k},\mathbf{v}_{k}):=(T_{k},\mathbf{w}_{k}).$
Then $\min_{i}(v_{k})_{i}=\min_{i}(w_{k})_{i}\to\infty$ by construction, and
$\frac{|S^{\prime}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}=\frac{|S\cap B_{k}|}{|B_{k}|}\longrightarrow\alpha,$
because $|S\cap B_{k}|/|B_{k}|>\alpha-\varepsilon_{k}-\varepsilon_{k}/100$ and $\varepsilon_{k}\to 0$. This completes the proof. ∎
The following two lemmas, Lemma 2.12 and Lemma 2.13, are used in the proof of Theorem 3.1. Lemma 2.12 provides a quantitative estimate for the proportion of uniformly KK–balanced vectors inside large boxes, while Lemma 2.13 ensures that any set of positive upper density contains a positive–density subset with bounded anisotropy. Together, they allow us to reduce the upper–bound argument to uniformly balanced regimes, where the covering and growth estimates apply effectively.
Lemma 2.12 (Density of $\mathcal{C}_{K}$ in $Q_{N}$).
Fix $d\geq 1$ and $K\geq 1$. Let
$Q_{N}:=[1,N]^{d}\cap\mathbb{N}^{d},\qquad\mathcal{C}_{K}:=\Bigl\{\mathbf{a}\in\mathbb{N}^{d}:\ \frac{\max_{j}a_{j}}{\min_{j}a_{j}}\leq K\Bigr\}.$
Then
$\lim_{N\to\infty}\frac{|\mathcal{C}_{K}\cap Q_{N}|}{|Q_{N}|}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1}.$
Equivalently,
$\lim_{N\to\infty}\frac{|Q_{N}\setminus\mathcal{C}_{K}|}{|Q_{N}|}=1-\Bigl(1-\frac{1}{K}\Bigr)^{d-1}.$
Proof.
Write $|Q_{N}|=N^{d}$ and set $M:=\lfloor N/K\rfloor$. For $\mathbf{a}=(a_{1},\dots,a_{d})\in Q_{N}$ let $m:=\min_{j}a_{j}$. The condition $\mathbf{a}\in\mathcal{C}_{K}$ is equivalent to $\max_{j}a_{j}\leq Km$. Hence, for each $m\in\{1,\dots,N\}$, all coordinates satisfy
$a_{j}\in[m,U_{m}]\cap\mathbb{N},\qquad U_{m}:=\min\{Km,N\}.$
Let $L_{m}:=|([m,U_{m}]\cap\mathbb{N})|=U_{m}-m+1$. The number of $\mathbf{a}\in Q_{N}$ with $\min_{j}a_{j}=m$ and $\max_{j}a_{j}\leq Km$ equals
$L_{m}^{d}-(L_{m}-1)^{d},$
since $L_{m}^{d}$ counts all dd–tuples in $m,U_{m}]^{d}$ and $(L_{m}-1)^{d}$ counts those with all coordinates at least $m+1$. Therefore,
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{N}\bigl(L_{m}^{d}-(L_{m}-1)^{d}\bigr).$ (6)
If $1\leq m\leq M$, then $Km\leq N$ so $U_{m}=Km$ and thus $L_{m}=(K-1)m+1$. If $M+1\leq m\leq N$, then $U_{m}=N$ and thus $L_{m}=N-m+1$. Splitting the sum in (6) gives
$|\mathcal{C}_{K}\cap Q_{N}|$ $=\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)+\sum_{m=M+1}^{N}\Bigl((N-m+1)^{d}-(N-m)^{d}\Bigr).$ (7)
The second sum telescopes: with t=Nm+1t=N-m+1 it becomes
$\sum_{t=1}^{N-M}\bigl(t^{d}-(t-1)^{d}\bigr)=(N-M)^{d}.$
Hence,
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)+(N-M)^{d}.$ (8)
For $x\geq 0$, the binomial expansion yields
$(x+1)^{d}-x^{d}=dx^{d-1}+O_{d}(x^{d-2})\qquad(d\geq 2),$
and for $d=1$ the statement is trivial. Applying this with x=(K1)mx=(K-1)m and summing over $m\leq M$ gives
$\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)=d(K-1)^{d-1}\sum_{m=1}^{M}m^{d-1}+O_{d,K}\Bigl(\sum_{m=1}^{M}m^{d-2}\Bigr).$
Using the standard power-sum estimates
$\sum_{m=1}^{M}m^{d-1}=\frac{M^{d}}{d}+O_{d}(M^{d-1}),\qquad\sum_{m=1}^{M}m^{d-2}=O_{d}(M^{d-1}),$
we obtain
$\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)=(K-1)^{d-1}M^{d}+O_{d,K}(M^{d-1}).$ (9)
Combining (8) and (9) yields
$|\mathcal{C}_{K}\cap Q_{N}|=(K-1)^{d-1}M^{d}+(N-M)^{d}+O_{d,K}(N^{d-1}).$
Dividing by $N^{d}$ and noting that $M/N\to 1/K$ as $N\to\infty$, we get
$\lim_{N\to\infty}\frac{|\mathcal{C}_{K}\cap Q_{N}|}{N^{d}}=(K-1)^{d-1}\Bigl(\frac{1}{K}\Bigr)^{d}+\Bigl(1-\frac{1}{K}\Bigr)^{d}.$
Finally,
$(K-1)^{d-1}\Bigl(\frac{1}{K}\Bigr)^{d}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1}\cdot\frac{1}{K},$
so the limit equals
$\Bigl(1-\frac{1}{K}\Bigr)^{d-1}\Bigl(\frac{1}{K}+1-\frac{1}{K}\Bigr)=\Bigl(1-\frac{1}{K}\Bigr)^{d-1},$
as claimed. ∎
Lemma 2.13 (Positive density forces bounded anisotropy).
Let $A\subset\mathbb{N}^{d}$ satisfy
$\overline{d}(A):=\limsup_{N\to\infty}\frac{|A\cap Q_{N}|}{|Q_{N}|}>0,\qquad Q_{N}=[1,N]^{d}\cap\mathbb{N}^{d}.$
Then there exists $K\geq 1$ such that
$\overline{d}\bigl(A\cap\mathcal{C}_{K}\bigr)>0,\qquad\mathcal{C}_{K}=\Bigl\{\mathbf{a}\in\mathbb{N}^{d}:\frac{\max_{j}a_{j}}{\min_{j}a_{j}}\leq K\Bigr\}.$
Proof.
Set $\delta:=\overline{d}(A)>0$. By Lemma 2.12 we have
$\lim_{N\to\infty}\frac{|\mathcal{C}_{K}\cap Q_{N}|}{|Q_{N}|}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1},$
and hence
$\lim_{N\to\infty}\frac{|Q_{N}\setminus\mathcal{C}_{K}|}{|Q_{N}|}=1-\Bigl(1-\frac{1}{K}\Bigr)^{d-1}.$
Choosing $K$ sufficiently large, we may ensure that
$1-\Bigl(1-\frac{1}{K}\Bigr)^{d-1}<\frac{\delta}{4}.$ (10)
By the definition of the limit, there exists $N_{1}$ such that for all $N\geq N_{1}$,
$\frac{|Q_{N}\setminus\mathcal{C}_{K}|}{|Q_{N}|}<\frac{\delta}{2}.$ (11)
By the definition of $\limsup$, there exists a sequence $N_{i}\to\infty$ such that
$\frac{|A\cap Q_{N_{i}}|}{|Q_{N_{i}}|}\longrightarrow\delta.$
In particular, for $i$ sufficiently large we have
$\frac{|A\cap Q_{N_{i}}|}{|Q_{N_{i}}|}>\frac{3\delta}{4},\qquad N_{i}\geq N_{1}.$ (12)
For such $i$, using the decomposition
$A\cap Q_{N_{i}}=\bigl(A\cap\mathcal{C}_{K}\cap Q_{N_{i}}\bigr)\;\cup\;\bigl(A\cap(Q_{N_{i}}\setminus\mathcal{C}_{K})\bigr),$
we obtain
$|A\cap\mathcal{C}_{K}\cap Q_{N_{i}}|\geq|A\cap Q_{N_{i}}|-|Q_{N_{i}}\setminus\mathcal{C}_{K}|.$
Dividing by $|Q_{N_{i}}|$ and combining (11) and (12), we conclude that
$\frac{|A\cap\mathcal{C}_{K}\cap Q_{N_{i}}|}{|Q_{N_{i}}|}\geq\frac{3\delta}{4}-\frac{\delta}{2}=\frac{\delta}{4}.$
Taking the $\limsup$ over $i$ yields
$\overline{d}\bigl(A\cap\mathcal{C}_{K}\bigr)\geq\frac{\delta}{4}>0,$
as claimed. ∎
2.3 Almost Lipschitz maps and elimination maps
We recall the notion of almost Lipschitz maps, which will be used to control the possible loss of Hausdorff dimension when digits are eliminated from symbolic addresses.
Definition 2.14.
Let $(X,d)$ be a compact metric space, $F\subset X$, and $f:F\to X$. We say that
  • $f$ is Hölder continuous with exponent $\gamma\in(0,1]$ if there exists $C>0$ such that
    $d\bigl(f(x),f(y)\bigr)\leq C\,d(x,y)^{\gamma}\quad\text{for all }x,y\in F;$
  • $f$ is almost Lipschitz if it is Hölder continuous with exponent $\gamma$ for every $\gamma\in(0,1)$.
The following dimension estimate is standard.
Lemma 2.15.
Let $(X,d)$ be a compact metric space, $F\subset X$, and let $f:F\to X$ be an almost Lipschitz map. Then
$\dim_{\mathrm{H}}f(F)\leq\dim_{\mathrm{H}}F.$
Proof.
See [7, Proposition 3.3]. ∎
In the sequel, we construct elimination maps which erase prescribed digits from symbolic addresses. Lemma 2.15 ensures that these operations do not increase the Hausdorff dimension.
3 Upper bounds on $\dim_{H}$ for digit–restricted sets in $\mathbb{N}^{d}$
Let $S\subset\mathbb{N}^{d}$. Recall that
$\widetilde{\mathcal{E}}_{S}:=\{\mathbf{x}\in(0,1)^{d}:\ \mathbf{a}_{n}(\mathbf{x})\in S\text{ for all }n\geq 1\},$
and its escaping part
$\mathcal{E}_{S}:=\Bigl\{\mathbf{x}\in\widetilde{\mathcal{E}}_{S}:\ a_{n}(x_{j})\to\infty\text{ for each }1\leq j\leq d\Bigr\}.$
We also consider the vector-injective subregime
$\mathcal{E}_{S}^{\mathrm{vec}}:=\Bigl\{\mathbf{x}\in\widetilde{\mathcal{E}}_{S}:\ \mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})\ (n\neq m),\ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\Bigr\}.$
Clearly, $\mathcal{E}_{S}^{\mathrm{vec}}\subset\mathcal{E}_{S}\subset\widetilde{\mathcal{E}}_{S}$. Hence any upper bound for $\dim_{H}(\widetilde{\mathcal{E}}_{S})$ yields an upper bound for $\dim_{H}(\mathcal{E}_{S})$ and $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})$, while any lower bound for $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})$ (or $\dim_{H}(\mathcal{E}_{S})$) also yields a lower bound for $\dim_{H}(\widetilde{\mathcal{E}}_{S})$.
3.1 Main upper bounds
We collect the main upper bounds proved in this section. Proofs are given in the subsections that follow.
Theorem 3.1.
If $S\subset\mathbb{N}^{d}$ has positive upper density, then
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=\dim_{H}(\mathcal{E}_{S})=\frac{d}{2}.$
Remark 3.2.
Assume that $S=S_{1}\times\cdots\times S_{d}\subset\mathbb{N}^{d}$. Then
$\mathcal{E}_{S}=\prod_{j=1}^{d}\mathcal{E}_{S_{j}},$
and the standard product inequalities give
$\sum_{j=1}^{d}\dim_{H}(\mathcal{E}_{S_{j}})\leq\dim_{H}(\mathcal{E}_{S})\leq\sum_{j=1}^{d}\dim_{P}(\mathcal{E}_{S_{j}}).$
In particular, if $\dim_{H}(\mathcal{E}_{S_{j}})=\alpha_{j}$ for all $j$, then $\dim_{H}(\mathcal{E}_{S})\geq\sum_{j=1}^{d}\alpha_{j}$, with equality provided $\dim_{P}(\mathcal{E}_{S_{j}})=\alpha_{j}$ for all $j$.
It is not clear in general whether $\dim_{P}(\mathcal{E}_{S_{j}})=\dim_{H}(\mathcal{E}_{S_{j}})$ holds for digit-restricted continued-fraction sets. Consequently, even in the product case $S=S_{1}\times\cdots\times S_{d}$, we do not know whether the additivity
$\dim_{H}(\mathcal{E}_{S})=\sum_{j=1}^{d}\alpha_{j}$
must hold under the sole assumption $\dim_{H}(\mathcal{E}_{S_{j}})=\alpha_{j}$.
Fix an integer $N\geq 2$. Define
$E_{\geq N}:=\bigl\{x\in(0,1):a_{n}(x)\geq N\text{ for all }n\geq 1\bigr\},\qquad\widetilde{\mathcal{E}}_{N}:=(E_{\geq N})^{d}\subset(0,1)^{d}.$
Theorem 3.3 (Asymptotics for the full restriction $a_{n}\geq N$).
For every $\varepsilon>0$ there exists $N_{0}(\varepsilon)$ such that for all $N\geq N_{0}(\varepsilon)$,
$\frac{d}{2}\leq\dim_{H}(\widetilde{\mathcal{E}}_{N})\leq\frac{d}{2}+\varepsilon.$
In particular, $\dim_{H}(\widetilde{\mathcal{E}}_{N})\to d/2$ as $N\to\infty$.
Theorem 3.4.
Let $S\subset\mathbb{N}^{d}$ be infinite. For $s\in(0,d]$ let $\phi^{s}$ be the singular value potential defined in (13) below, and set
$\Phi_{S}(s):=\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a}),\qquad s_{\sharp}:=\inf\{s>0:\Phi_{S}(s)<1\},\qquad s_{\ast}:=\inf\{s>0:\Phi_{S}(s)<\infty\}.$
Define also the multivariate Dirichlet series
$\zeta_{S}(\bm{\sigma}):=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\sigma_{j}}\quad(\bm{\sigma}\in(0,\infty)^{d}),$
its convergence domain $\mathcal{D}_{S}:=\{\bm{\sigma}:\zeta_{S}(\bm{\sigma})<\infty\}$, and the trace exponent
$\Lambda_{S}:=\inf_{\bm{\sigma}\in\mathcal{D}_{S}}(\sigma_{1}+\cdots+\sigma_{d}).$
Then:
  1. (1)
    $\dim_{H}(\widetilde{\mathcal{E}}_{S})\leq s_{\sharp}$.
  2. (2)
    $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}({\mathcal{E}}_{S})\leq s_{\ast}$
  3. (3)
    $\Lambda_{S}\leq 2s_{\ast}$ (hence $\tfrac{1}{2}\Lambda_{S}\leq s_{\ast}\leq s_{\sharp}$).
  4. (4)
    If $S$ is uniformly KK–balanced (i.e. $\max_{i}a_{i}\leq K\min_{i}a_{i}$ for all $\mathbf{a}\in S$), then $s_{\ast}=\tfrac{1}{2}\Lambda_{S}$.
Corollary 3.5 (Growth bound in the balanced region).
Let $K\geq 1$ and let $S\subset\mathcal{C}_{K}\subset\mathbb{N}^{d}$ be infinite. Assume that there exists $\alpha\geq 1$ such that
$|S\cap Q_{N}|\ \ll\ N^{d/\alpha}\qquad(N\to\infty),\qquad Q_{N}:=[1,N]^{d}\cap\mathbb{N}^{d}.$
Then
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\ \leq\ \frac{d}{2\alpha}.$
Proof of Corollary 3.5 assuming Theorem 3.4:.
Fix $\varepsilon>0$ and set $\sigma:=\frac{1}{\alpha}+\varepsilon$ and $\bm{\sigma}:=(\sigma,\dots,\sigma)$. We show that $\bm{\sigma}\in\mathcal{D}_{S}$, i.e. $\zeta_{S}(\bm{\sigma})<\infty$.
Decompose $S$ into dyadic shells $Q_{2^{n+1}}\setminus Q_{2^{n}}$. Since $|S\cap Q_{2^{n+1}}|\ll 2^{(n+1)d/\alpha}$ and $S\subset\mathcal{C}_{K}$, for every $\mathbf{a}\in S\cap(Q_{2^{n+1}}\setminus Q_{2^{n}})$ we have $\max_{j}a_{j}>2^{n}$ and hence $\min_{j}a_{j}\geq K^{-1}\max_{j}a_{j}>2^{n}/K$, so
$\prod_{j=1}^{d}a_{j}^{-\sigma}\leq\Bigl(\frac{2^{n}}{K}\Bigr)^{-d\sigma}.$
Therefore,
$\zeta_{S}(\bm{\sigma})\leq\sum_{n\geq 0}|S\cap Q_{2^{n+1}}|\cdot\Bigl(\frac{2^{n}}{K}\Bigr)^{-d\sigma}\ll_{d,\alpha,K,\sigma}\sum_{n\geq 0}2^{nd(1/\alpha-\sigma)}<\infty,$
since $\sigma>1/\alpha$. Hence $\bm{\sigma}\in\mathcal{D}_{S}$ and
$\Lambda_{S}\leq\sigma_{1}+\cdots+\sigma_{d}=d\sigma=\frac{d}{\alpha}+d\varepsilon.$
Because $S$ is uniformly KK–balanced, Theorem 3.4(4) gives $s_{\ast}=\tfrac{1}{2}\Lambda_{S}$, so
$s_{\ast}\leq\frac{d}{2\alpha}+\frac{d\varepsilon}{2}.$
Finally, by Theorem 3.4(2), we have
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}\leq\frac{d}{2\alpha}+\frac{d\varepsilon}{2}.$
Letting $\varepsilon\downarrow 0$ yields $\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\leq\frac{d}{2\alpha}$. ∎
3.2 Universal upper bounds via singular values and Dirichlet series
In this subsection we develop a collection of general tools for obtaining upper bounds on the Hausdorff dimension of digit–restricted sets. We show that finite modifications of digit constraints do not affect dimension, establish a universal upper bound in terms of a singular–value type potential, and relate this geometric quantity to a multivariate Dirichlet series invariant.
Lemma 3.6 (Digit insertion is bi-Lipschitz).
Let $X\subset(0,1)^{d}$ and fix digit vectors $\mathbf{a}_{1},\dots,\mathbf{a}_{N}\in\mathbb{N}^{d}$. Define $\Phi:X\to(0,1)^{d}$ by inserting $\mathbf{a}_{1},\dots,\mathbf{a}_{N}$ as the first $N$ digits in each coordinate, i.e. for $\mathbf{y}=(y_{1},\dots,y_{d})\in X$,
$x_{j}=\frac{1}{a_{1,j}+\displaystyle\frac{1}{a_{2,j}+\cdots+\displaystyle\frac{1}{a_{N,j}+y_{j}}}},\qquad 1\leq j\leq d.$
Let $X_{\mathbf{a}_{1},\dots,\mathbf{a}_{N}}:=\Phi(X)$. Then
$\dim_{H}(X_{\mathbf{a}_{1},\dots,\mathbf{a}_{N}})=\dim_{H}(X).$
Proof.
For each coordinate, the map $y_{j}\mapsto x_{j}$ is a Möbius transformation
$x_{j}=\frac{P_{N,j}+y_{j}P_{N-1,j}}{Q_{N,j}+y_{j}Q_{N-1,j}},$
with convergents determined by $a_{1,j},\dots,a_{N,j}$. On $y_{j}\in(0,1)$ the derivative is bounded above and below by positive constants depending only on the prescribed digits. Hence each coordinate map is bi-Lipschitz on $(0,1)$ with uniform constants, and the product map $\Phi$ is bi-Lipschitz on $(0,1)^{d}$. Hausdorff dimension is invariant under bi-Lipschitz maps. ∎
Lemma 3.7 (Deleting finitely many digits does not change $\dim_{H}$).
Fix $S\subset\mathbb{N}^{d}$ and $N\in\mathbb{N}$. Then
$\dim_{H}(\mathcal{E}_{S})=\dim_{H}\Bigl\{\mathbf{x}:\ \mathbf{a}_{n}(\mathbf{x})\in S\ (n\geq N),\ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\Bigr\}.$
The same statement holds with $\mathcal{E}_{S}$ replaced by $\widetilde{\mathcal{E}}_{S}$.
Proof.
Apply Lemma 3.6 to insert/delete the first N1N-1 digit vectors. The relevant set is a countable union over all choices of the first N1N-1 digits, and each branch is bi-Lipschitz equivalent to the tail set. ∎
For $\mathbf{a}=(a_{1},\dots,a_{d})\in\mathbb{N}^{d}$ let $a_{(1)}\leq\cdots\leq a_{(d)}$ be the nondecreasing rearrangement of its coordinates. For $s\in(0,d]$ let $k$ be the unique integer with $k<s\leq k+1$ and define
$\phi^{s}(\mathbf{a}):=\Bigl(\prod_{j=1}^{k}a_{(j)}^{-2}\Bigr)\cdot a_{(k+1)}^{-2(s-k)}.$ (13)
For a family $(S_{n})_{n\geq 1}$ with $S_{n}\subset\mathbb{N}^{d}$ define the corresponding sets
$\mathcal{E}_{(S_{n})}:=\Bigl\{\mathbf{x}\in(0,1)^{d}:\mathbf{a}_{n}(\mathbf{x})\in S_{n}\ (n\geq 1),\ \ a_{n}(x_{j})\to\infty\ (1\leq j\leq d)\Bigr\},$
and similarly $\widetilde{\mathcal{E}}_{(S_{n})}$ without the divergence requirement.
Proposition 3.8 (Universal upper bound).
Let $(S_{n})_{n\geq 1}$ be subsets of $\mathbb{N}^{d}$ and fix $s\in(0,d]$. Set
$\Phi_{S_{n}}(s):=\sum_{\mathbf{a}\in S_{n}}\phi^{s}(\mathbf{a}).$
If
$\limsup_{n\to\infty}\Bigl(\prod_{k=1}^{n}\Phi_{S_{k}}(s)\Bigr)^{1/n}<1,$
then $\dim_{H}(\widetilde{\mathcal{E}}_{(S_{n})})\leq s$, and hence also $\dim_{H}(\mathcal{E}_{(S_{n})})\leq s$. In particular, for fixed $S$ one has $\dim_{H}(\widetilde{\mathcal{E}}_{S})\leq s$ whenever $\Phi_{S}(s)<1$.
Proof.
Fix $s\in(0,d]$ and let $k\in\{0,1,\dots,d-1\}$ be the unique integer such that $k<s\leq k+1$. Assume that
$\limsup_{n\to\infty}\Bigl(\prod_{j=1}^{n}\Phi_{S_{j}}(s)\Bigr)^{1/n}<1.$
Then there exist $\rho\in(0,1)$ and $n_{0}\in\mathbb{N}$ such that
$\prod_{j=1}^{n}\Phi_{S_{j}}(s)\leq\rho^{n}\qquad\text{for all }n\geq n_{0}.$ (14)
For each $n\geq 1$ we have the cylinder cover
$\widetilde{\mathcal{E}}_{(S_{n})}\subset\bigcup_{\omega_{n}\in S_{1}\times\cdots\times S_{n}}I_{\omega_{n}},$ (15)
where $\omega_{n}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ and
$I_{\omega_{n}}=\prod_{j=1}^{d}I(a_{1,j},\dots,a_{n,j})$
is the product cylinder. Write $L_{j}(\omega_{n})$ for the side length of $I_{\omega_{n}}$ in coordinate $j$. By Lemma 2.4 and Lemma 2.1(i), there exist constants $c_{0},C_{0}>0$ such that for every $1\leq j\leq d$,
$c_{0}\prod_{i=1}^{n}a_{i,j}^{-2}\ \leq\ L_{j}(\omega_{n})\ \leq\ C_{0}\prod_{i=1}^{n}a_{i,j}^{-2}.$ (16)
Let $\mathcal{L}_{1}(\omega_{n})\geq\cdots\geq\mathcal{L}_{d}(\omega_{n})$ be the nonincreasing rearrangement of the side lengths $\{L_{j}(\omega_{n})\}_{j=1}^{d}$ and set
$\varphi^{s}(I_{\omega_{n}}):=\mathcal{L}_{1}(\omega_{n})\cdots\mathcal{L}_{k}(\omega_{n})\,\mathcal{L}_{k+1}(\omega_{n})^{\,s-k}.$ (17)
Recall that the ss–dimensional Hausdorff content of a set $E$ is
$\mathcal{H}^{s}_{\infty}(E)=\inf\Bigl\{\sum_{m}\operatorname{diam}(U_{m})^{s}:\ E\subset\bigcup_{m}U_{m}\Bigr\}.$
A standard covering argument for axis-parallel rectangles yields a constant $C_{1}=C_{1}(d,s)>0$ such that, for every $\omega_{n}$,
$\mathcal{H}^{s}_{\infty}(I_{\omega_{n}})\leq C_{1}\,\varphi^{s}(I_{\omega_{n}}).$ (18)
Using (16) and the definition (13), there is a constant $C_{2}=C_{2}(d,s)>0$ such that for every word $\omega_{n}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$,
$\varphi^{s}(I_{\omega_{n}})\leq C_{2}\prod_{i=1}^{n}\phi^{s}(\mathbf{a}_{i}).$ (19)
Define
$\mathcal{W}_{n}(s):=\sum_{\omega_{n}\in S_{1}\times\cdots\times S_{n}}\varphi^{s}(I_{\omega_{n}}).$
Then (19) gives
$\mathcal{W}_{n}(s)\leq C_{2}\sum_{\omega_{n}\in S_{1}\times\cdots\times S_{n}}\ \prod_{i=1}^{n}\phi^{s}(\mathbf{a}_{i})=C_{2}\prod_{i=1}^{n}\Bigl(\sum_{\mathbf{a}\in S_{i}}\phi^{s}(\mathbf{a})\Bigr)=C_{2}\prod_{i=1}^{n}\Phi_{S_{i}}(s).$
Combining this with (15) and (18), we obtain for all $n\geq 1$,
$\mathcal{H}^{s}_{\infty}\bigl(\widetilde{\mathcal{E}}_{(S_{n})}\bigr)\leq\sum_{\omega_{n}\in S_{1}\times\cdots\times S_{n}}\mathcal{H}^{s}_{\infty}(I_{\omega_{n}})\leq C_{1}\,\mathcal{W}_{n}(s)\leq C_{1}C_{2}\prod_{i=1}^{n}\Phi_{S_{i}}(s).$
By (14), the right-hand side tends to 0 as $n\to\infty$. Hence $\mathcal{H}^{s}_{\infty}(\widetilde{\mathcal{E}}_{(S_{n})})=0$ and therefore $\dim_{H}(\widetilde{\mathcal{E}}_{(S_{n})})\leq s$. Since $\mathcal{E}_{(S_{n})}\subset\widetilde{\mathcal{E}}_{(S_{n})}$, the same upper bound holds for $\mathcal{E}_{(S_{n})}$.
In the stationary case $S_{n}\equiv S$, the hypothesis reduces to $\Phi_{S}(s)<1$, and the conclusion follows immediately. ∎
Define
$\zeta_{S}(\bm{\sigma}):=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\sigma_{j}},\qquad\mathcal{D}_{S}:=\{\bm{\sigma}\in(0,\infty)^{d}:\ \zeta_{S}(\bm{\sigma})<\infty\},$
and the trace infimum
$\Lambda_{S}:=\inf_{\bm{\sigma}\in\mathcal{D}_{S}}(\sigma_{1}+\cdots+\sigma_{d}).$
We will need a small combinatorial device to compare weighted sums after sorting coordinates; for completeness we record the following rearrangement inequality.
Lemma 3.9 (Rearrangement inequality).
Let $x_{1}\leq\cdots\leq x_{d}$ and $w_{1}\geq\cdots\geq w_{d}$ be real numbers. Then for every permutation $\pi\in S_{d}$,
$\sum_{i=1}^{d}w_{\pi(i)}x_{i}\ \geq\ \sum_{i=1}^{d}w_{i}x_{i}.$ (20)
Proof.
If $p<q$ and $w_{\pi(p)}<w_{\pi(q)}$, let $\pi^{\prime}$ be obtained from $\pi$ by swapping $\pi(p)$ and $\pi(q)$. Then
$\sum_{i=1}^{d}w_{\pi(i)}x_{i}-\sum_{i=1}^{d}w_{\pi^{\prime}(i)}x_{i}=(w_{\pi(q)}-w_{\pi(p)})(x_{q}-x_{p})\geq 0.$
Iterating this swap removes inversions and yields (20). ∎
Proposition 3.10 (Zeta trace controls $\Phi_{S}$).
Let $S\subset\mathbb{N}^{d}$ be infinite and define
$s_{\ast}:=\inf\{s>0:\ \Phi_{S}(s)<\infty\}.$
Then
$\Lambda_{S}\leq 2s_{\ast}.$
Proof.
Fix $s>s_{\ast}$. Then $\Phi_{S}(s)=\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty$. Let $k\in\{0,1,\dots,d-1\}$ be the unique integer such that $k<s\leq k+1$.
Define the exponent vector
$\widetilde{\bm{\sigma}}(s):=(\underbrace{2,\dots,2}_{k\ \text{times}},\ 2(s-k),\ \underbrace{0,\dots,0}_{d-k-1\ \text{times}})\in[0,\infty)^{d},$
whose coordinates are non-increasing. For $\mathbf{a}=(a_{1},\dots,a_{d})\in\mathbb{N}^{d}$, let $a_{(1)}\leq\cdots\leq a_{(d)}$ be the nondecreasing rearrangement of its coordinates and set $x_{i}:=\log a_{(i)}$, so $x_{1}\leq\cdots\leq x_{d}$.
Let $\tau\in S_{d}$ be the permutation such that $a_{\tau(i)}=a_{(i)}$ for $1\leq i\leq d$. Applying Lemma 3.9 with the nondecreasing sequence $(x_{i})_{i=1}^{d}$ and the nonincreasing weights $(\widetilde{\sigma}_{i}(s))_{i=1}^{d}$, and taking $\pi=\tau$, we obtain
$\sum_{i=1}^{d}\widetilde{\sigma}_{\tau(i)}(s)\,x_{i}\ \geq\ \sum_{i=1}^{d}\widetilde{\sigma}_{i}(s)\,x_{i}=2\sum_{i=1}^{k}x_{i}+2(s-k)x_{k+1}.$
Since $\sum_{i=1}^{d}\widetilde{\sigma}_{\tau(i)}(s)\,x_{i}=\sum_{j=1}^{d}\widetilde{\sigma}_{j}(s)\,\log a_{j}$, this yields
$\sum_{j=1}^{d}\widetilde{\sigma}_{j}(s)\,\log a_{j}\ \geq\ 2\sum_{i=1}^{k}\log a_{(i)}+2(s-k)\log a_{(k+1)}.$
Exponentiating and using the definition of $\phi^{s}$ gives the pointwise bound
$\prod_{j=1}^{d}a_{j}^{-\widetilde{\sigma}_{j}(s)}\leq\Bigl(\prod_{i=1}^{k}a_{(i)}^{-2}\Bigr)\,a_{(k+1)}^{-2(s-k)}=\phi^{s}(\mathbf{a}).$
Summing over $\mathbf{a}\in S$ gives
$\zeta_{S}\bigl(\widetilde{\bm{\sigma}}(s)\bigr)=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\widetilde{\sigma}_{j}(s)}\leq\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})=\Phi_{S}(s)<\infty.$
Hence $\widetilde{\bm{\sigma}}(s)\in\mathcal{D}_{S}$ and therefore
$\Lambda_{S}\leq\sum_{j=1}^{d}\widetilde{\sigma}_{j}(s)=2s.$
Letting $s\downarrow s_{\ast}$ completes the proof. ∎
Remark 3.11.
The convergence region $\mathcal{D}_{S}$ is a global summability invariant and may be dominated by a single thick coordinate fiber. In particular, even if $S$ is sparse in most directions, the presence of one coordinate fiber with slow decay can force $\Lambda_{S}$ to be large. This reflects the fact that $\zeta_{S}$ records worst–case behavior across coordinates rather than any frequency or distributional information.
If $S$ is uniformly KK–balanced, then the above pathology disappears. Indeed, for $\sigma>0$ one has
$\prod_{j=1}^{d}a_{j}^{-\sigma}\asymp_{K}a_{(1)}^{-\sigma d},\qquad\phi^{s}(\mathbf{a})\asymp_{K}\prod_{j=1}^{d}a_{j}^{-2s/d}.$
Consequently, the diagonal Dirichlet series $\zeta_{S}(\sigma,\dots,\sigma)$ and the scalar series $\Phi_{S}(s)$ become equivalent after the change of variables $\sigma=2s/d$, and in this case
$s_{\ast}=\tfrac{1}{2}\,\Lambda_{S}.$
3.3 Proof of Theorem 3.4
Proof of Theorem 3.4.
(1) By Proposition 3.8 with $S_{n}\equiv S$, if $\Phi_{S}(s)<1$ then $\dim_{H}(\widetilde{\mathcal{E}}_{S})\leq s$. Taking the infimum over such $s$ yields $\dim_{H}(\widetilde{\mathcal{E}}_{S})\leq s_{\sharp}$.
(2) Fix any $s>s_{\ast}$. Then $\Phi_{S}(s)=\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty$. For $M\in\mathbb{N}$ set
$S^{\geq M}:=\{\mathbf{a}\in S:\ a_{(1)}(\mathbf{a})=\min_{j}a_{j}\geq M\}.$
Since $(S^{\geq M})_{M\geq 1}$ is a decreasing family with $\bigcap_{M\geq 1}S^{\geq M}=\varnothing$, monotone convergence gives $\Phi_{S^{\geq M}}(s)\downarrow 0$ as $M\to\infty$. Hence we may choose $M$ such that
$\Phi_{S^{\geq M}}(s)<1.$ (21)
Now let $\mathcal{E}_{S}^{(M,N)}$ be the set of points $\mathbf{x}\in(0,1)^{d}$ such that $\mathbf{a}_{n}(\mathbf{x})\in S$ for $1\leq n<N$, $\mathbf{a}_{n}(\mathbf{x})\in S^{\geq M}$ for all $n\geq N$, and $a_{n}(x_{j})\to\infty$ for each $1\leq j\leq d$. Since $\min_{j}a_{n}(x_{j})\to\infty$ for $\mathbf{x}\in\mathcal{E}_{S}$, every $\mathbf{x}\in\mathcal{E}_{S}$ belongs to $\mathcal{E}_{S}^{(M,N)}$ for some $N$, and therefore
$\mathcal{E}_{S}\subset\bigcup_{N\geq 1}\mathcal{E}_{S}^{(M,N)}.$
By Lemma 3.7, deleting (or inserting) finitely many initial digit vectors does not change Hausdorff dimension, hence for each $N$,
$\dim_{H}(\mathcal{E}_{S}^{(M,N)})=\dim_{H}(\mathcal{E}_{S^{\geq M}}).$
Consequently, since the Hausdorff dimension of a countable union is the supremum of the Hausdorff dimensions of the pieces, we obtain
$\dim_{H}(\mathcal{E}_{S})\leq\dim_{H}\Bigl(\bigcup_{N\geq 1}\mathcal{E}_{S}^{(M,N)}\Bigr)\leq\sup_{N\geq 1}\dim_{H}(\mathcal{E}_{S}^{(M,N)})=\dim_{H}(\mathcal{E}_{S^{\geq M}}).$
Finally, applying Proposition 3.8 to the stationary family $S_{n}\equiv S^{\geq M}$ and using (21), we obtain $\dim_{H}(\mathcal{E}_{S^{\geq M}})\leq s$, and therefore $\dim_{H}(\mathcal{E}_{S})\leq s$. Since $s>s_{\ast}$ was arbitrary, it follows that $\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}$. In particular,
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}.$
(3) This is Proposition 3.10.
(4) Suppose that $S$ is uniformly KK–balanced, that is, $\max_{j}a_{j}\leq K\min_{j}a_{j}$ for all $\mathbf{a}\in S$. Then for every $\sigma>0$ and $\mathbf{a}\in S$,
$\prod_{j=1}^{d}a_{j}^{-\sigma}\asymp_{K}a_{(1)}^{-\sigma d},$
where $a_{(1)}=\min_{j}a_{j}$. Similarly, for $s\in(0,d]$ one has
$\phi^{s}(\mathbf{a})=\Bigl(\prod_{j=1}^{k}a_{(j)}^{-2}\Bigr)\,a_{(k+1)}^{-2(s-k)}\asymp_{K}\prod_{j=1}^{d}a_{j}^{-2s/d},$
with implicit constants depending only on $K,d,s$.
Consequently, after the change of variables $\sigma=2s/d$, the diagonal Dirichlet series and the singular series are equivalent in the sense that
$\zeta_{S}(\sigma,\dots,\sigma)<\infty\quad\Longleftrightarrow\quad\Phi_{S}(s)<\infty.$
By definition of $\Lambda_{S}$ and $s_{\ast}$, this implies
$\Lambda_{S}=2s_{\ast},$
and hence $s_{\ast}=\tfrac{1}{2}\,\Lambda_{S}$ in the uniformly KK–balanced case.
3.4 Proof of Theorem 3.3
A key step is to bound the Hausdorff dimension of certain product sets arising from one–dimensional tail continued–fraction sets. We reduce the dd–dimensional coordinatewise problem to one dimension using two standard inputs: (i) for tail alphabets in dimension one, $\dim_{H}=\dim_{B}=\dim_{P}$; and (ii) a product inequality for Hausdorff dimension whose upper bound involves packing dimension. We will also use a quantitative estimate for $\dim_{H}(E_{\geq N})$ due to Good.
For $N\geq 2$ set
$E_{\geq N}:=\{x\in(0,1):a_{n}(x)\geq N\text{ for all }n\geq 1\}.$
Lemma 3.12.
We have
$\dim_{H}(E_{\geq N})=\dim_{B}(E_{\geq N})=\dim_{P}(E_{\geq N}),$
where $\dim_{B}$ and $\dim_{P}$ denote the (upper) box and packing dimensions.
Proof.
This is a standard fact for one–dimensional continued fractions with a tail alphabet. A direct reference is [22, Section 6, Remark, p. 5023], which records the coincidence of Hausdorff, box, and packing dimensions in this setting.
Alternatively, one may combine [22, Corollary 5.9] with the general identity $\dim_{P}=\dim_{B}$ for conformal iterated function systems ([21, Theorem 3.1]) and the discussion in [22, Section 6]. ∎
Remark 3.13.
We emphasize that the coincidence in Lemma 3.12 is not automatic for general conformal iterated function systems. In fact, Mauldin–Urbański [22, Theorem 6.2] exhibited many examples where Hausdorff dimension is strictly smaller than box/packing dimension: for every $p\geq 2$ there exists $q\geq 1$ such that if $l\geq q$ and $I_{l}:=\{n^{p}:n\geq l\}$, then the corresponding continued–fraction digit–restricted set satisfies
$\dim_{H}(\widetilde{\mathcal{E}}_{I_{l}})<\underline{\dim}_{B}(\widetilde{\mathcal{E}}_{I_{l}})\leq\overline{\dim}_{B}(\widetilde{\mathcal{E}}_{I_{l}})=\dim_{P}(\widetilde{\mathcal{E}}_{I_{l}}).$
Here $\underline{\dim}_{B}$ and $\overline{\dim}_{B}$ denote the lower and upper box dimensions, respectively, and $\dim_{P}$ denotes the packing dimension. In particular, $\dim_{H}=\dim_{B}=\dim_{P}$ requires additional structure (such as the tail–alphabet situation) and should not be viewed as a general feature of conformal IFS.
We will use the following product inequality; see, e.g., [19].
Lemma 3.14.
Suppose that $A\subset\mathbb{R}^{d}$ and $B\subset\mathbb{R}^{t}$ are Borel sets. Then
$\dim_{H}A+\dim_{H}B\ \leq\ \dim_{H}(A\times B)\ \leq\ \dim_{H}A+\dim_{P}B,$ (22)
where $\dim_{P}$ denotes the packing dimension.
Lemma 3.15 (Good’s estimate for large partial quotients [10]).
Assume $N\geq 20$. Then
$\frac{1}{2}+\frac{1}{2\log(N+2)}\;<\;\dim_{H}(E_{\geq N})\;<\;\frac{1}{2}+\frac{\log\log(N-1)}{2\log(N-1)}.$
Proof of Theorem $3.$
By [10, Theorem 1], the set
$E_{\infty}:=\{x\in(0,1):a_{n}(x)\to\infty\}$
satisfies $\dim_{H}(E_{\infty})=\tfrac{1}{2}$. Since $\mathcal{E}_{\mathbb{N}^{d}}=E_{\infty}^{d}$, iterating the left inequality in Lemma 3.14 yields
$\dim_{H}(\mathcal{E}_{\mathbb{N}^{d}})=\dim_{H}(E_{\infty}^{d})\geq d\,\dim_{H}(E_{\infty})=\frac{d}{2}.$
Moreover, $\mathcal{E}_{\mathbb{N}^{d}}\subset\widetilde{\mathcal{E}}_{N}$ for every $N$, and therefore
$\frac{d}{2}\leq\dim_{H}(\widetilde{\mathcal{E}}_{N})\qquad\text{for all }N.$ (23)
Next, for $N\geq 2$ set
$E_{\geq N}:=\{x\in(0,1):a_{n}(x)\geq N\text{ for all }n\geq 1\},\qquad\widetilde{\mathcal{E}}_{N}=(E_{\geq N})^{d}.$
By Lemma 3.12 we have
$\dim_{H}(E_{\geq N})=\dim_{B}(E_{\geq N})=\dim_{P}(E_{\geq N}).$ (24)
We first obtain an upper bound for $\dim_{H}(\widetilde{\mathcal{E}}_{N})$ in terms of $\dim_{H}(E_{\geq N})$. By Lemma 3.14 and (24),
$\dim_{H}\bigl((E_{\geq N})\times(E_{\geq N})\bigr)\leq\dim_{H}(E_{\geq N})+\dim_{P}(E_{\geq N})=2\,\dim_{H}(E_{\geq N}).$
Iterating this estimate gives
$\dim_{H}(\widetilde{\mathcal{E}}_{N})=\dim_{H}\bigl((E_{\geq N})^{d}\bigr)\leq d\,\dim_{H}(E_{\geq N}).$ (25)
On the other hand, Lemma 3.15 implies that $\dim_{H}(E_{\geq N})\to\tfrac{1}{2}$ as $N\to\infty$. Combining this with (25) yields
$\limsup_{N\to\infty}\dim_{H}(\widetilde{\mathcal{E}}_{N})\leq\limsup_{N\to\infty}d\,\dim_{H}(E_{\geq N})=\frac{d}{2}.$
Together with (23), we conclude that
$\lim_{N\to\infty}\dim_{H}(\widetilde{\mathcal{E}}_{N})=\frac{d}{2}.$
4 Moran Fractal Constructions in Higher Dimensions
This section recalls a standard lower bound for Moran–type constructions in $\mathbb{R}^{d}$ using the mass distribution principle; see [7, 20] for background. The resulting bounds need not be sharp in general, and all constructions are based on coordinate cubes.
Definition 4.1 (Moran fractal).
Let $d\in\mathbb{N}$. Let $(r_{n})_{n\geq 1}$ and $(\delta_{n})_{n\geq 1}$ satisfy $r_{n}\geq 2$ and $0<\delta_{n+1}<\delta_{n}$ for all $n$. A compact set $F\subset[0,1]^{d}$ is called a Moran fractal with parameters $(r_{n})$ and $(\delta_{n})$ if there exist families $\mathcal{Q}_{n}$ of finitely many pairwise disjoint closed cubes in $0,1]^{d}$ such that, writing
$E_{n}:=\bigcup_{Q\in\mathcal{Q}_{n}}Q,\qquad F:=\bigcap_{n\geq 0}E_{n},$
the following hold for every $n\geq 1$:
  1. 1.
    (Nestedness) For each $Q\in\mathcal{Q}_{n}$ there is a unique parent $\widehat{Q}\in\mathcal{Q}_{n-1}$ with $Q\subset\widehat{Q}$.
  2. 2.
    (Branching) Each parent cube $P\in\mathcal{Q}_{n-1}$ contains at least $r_{n}$ distinct children from $\mathcal{Q}_{n}$.
  3. 3.
    (Separation inside a parent) If $Q,Q^{\prime}\in\mathcal{Q}_{n}$ are distinct and have the same parent in $\mathcal{Q}_{n-1}$, then $\operatorname{dist}(Q,Q^{\prime})\geq\delta_{n}$.
  4. 4.
    (Vanishing mesh) $\max_{Q\in\mathcal{Q}_{n}}\operatorname{diam}(Q)\to 0$ as $n\to\infty$.
Lemma 4.2 (Generalized Moran’s formula, lower bound).
Let $F\subset[0,1]^{d}$ be a Moran fractal with parameters $(r_{n})$ and $(\delta_{n})$. Then
$\dim_{H}F\;\geq\;\liminf_{n\to\infty}\frac{\log(r_{1}\cdots r_{n-1})}{-\log(r_{n}^{1/d}\delta_{n})}.$ (26)
The proof of Lemma 4.2 relies on several lemmas. We state them first.
Lemma 4.3 (Packing estimate in $\mathbb{R}^{d}$).
Let $d\in\mathbb{N}$. There exists a constant $C_{d}\geq 1$ depending only on $d$ such that the following holds.
Let $A\subset\mathbb{R}^{d}$ be bounded and let $\delta>0$. Suppose $S\subset A$ is finite and
$\|x-y\|\geq\delta\qquad\text{for all distinct }x,y\in S.$
Then
$S\leq C_{d}\Bigl(\frac{\operatorname{diam}(A)}{\delta}+1\Bigr)^{d}.$
In particular, if $S\geq 2$, then
$\operatorname{diam}(A)\ \geq\ c_{d}\,(\#S)^{1/d}\,\delta$
for some constant $c_{d}>0$ depending only on $d$.
Proof.
Write m:=#Sm:=\#S and $D:=\operatorname{diam}(A)$. Fix $x_{0}\in A$. Then $A\subset B(x_{0},D)$. For each $x\in S$ consider the closed ball $B(x,\delta/2)$. These balls are pairwise disjoint since $\|x-y\|\geq\delta$ for $x\neq y$. Moreover, for any $y\in B(x,\delta/2)$,
$\|y-x_{0}\|\leq\|y-x\|+\|x-x_{0}\|\leq\frac{\delta}{2}+D,$
so $B(x,\delta/2)\subset B(x_{0},D+\delta/2)$.
Let $\operatorname{vol}\bigl(B(0,1)\bigr)$ denote the volume of the unit ball in $\mathbb{R}^{d}$. Comparing volumes gives
$m\cdot\operatorname{vol}\bigl(B(0,1)\bigr)\Bigl(\frac{\delta}{2}\Bigr)^{d}\leq\operatorname{vol}\Bigl(B\Bigl(x_{0},D+\frac{\delta}{2}\Bigr)\Bigr)=\operatorname{vol}\bigl(B(0,1)\bigr)\Bigl(D+\frac{\delta}{2}\Bigr)^{d}.$
Cancelling $v_{d}$ yields
$m\leq\Bigl(\frac{D+\delta/2}{\delta/2}\Bigr)^{d}=\Bigl(\frac{2D}{\delta}+1\Bigr)^{d}\leq 3^{d}\Bigl(\frac{D}{\delta}+1\Bigr)^{d},$
which proves the first claim with $C_{d}:=3^{d}$.
For the second claim, assume $m\geq 2$. Then $D\geq\delta$ (since $S$ contains two points at distance at least $\delta$), hence $2D/\delta+1\leq 3D/\delta$. Therefore
$m\leq\Bigl(\frac{2D}{\delta}+1\Bigr)^{d}\leq\Bigl(\frac{3D}{\delta}\Bigr)^{d},$
so $D\geq 3^{-1}m^{1/d}\delta$. Thus we may take $c_{d}:=1/3$. ∎
We now state the second lemma, which establishes an upper bound for the lower limit.
Lemma 4.4.
With the notation of Lemma 4.2, assume moreover that $r_{k}\geq 2$ for all $k\in\mathbb{N}$. Then
$\liminf_{n\to\infty}\frac{\log(r_{1}\cdots r_{n-1})}{-\log L_{n}}\;\leq\;d,$
where $L_{n}:=r_{n}^{1/d}\delta_{n}$. In particular, $L_{n}\to 0$ as $n\to\infty$, and therefore the right-hand side of (26) in Lemma 4.2 is well defined (for all sufficiently large $n$ one has $0<L_{n}<1$ and $\log L_{n}>0$).
Proof.
Write $N_{n}:=\#\mathcal{E}_{n}$. Since each (k1)(k-1)-cube has at least $r_{k}$ children,
$N_{n}\geq r_{1}\cdots r_{n}\qquad(n\geq 1).$
We claim that $\mathcal{E}_{n}$ is $\delta_{n}$–separated. Take $Q\neq Q^{\prime}\in\mathcal{E}_{n}$. For $0\leq k\leq n$ let $Q^{(k)}$ and $(Q^{\prime})^{(k)}$ be the unique $k$-level ancestors. Let $m$ be the least index with $Q^{(m)}\neq(Q^{\prime})^{(m)}$. Then $Q^{(m-1)}=(Q^{\prime})^{(m-1)}$, so $Q^{(m)}$ and $(Q^{\prime})^{(m)}$ are two distinct children of the same parent. Hence
$\operatorname{dist}\bigl(Q^{(m)},(Q^{\prime})^{(m)}\bigr)\geq\delta_{m}.$
Since $Q\subset Q^{(m)}$ and $Q^{\prime}\subset(Q^{\prime})^{(m)}$, we have
$\operatorname{dist}(Q,Q^{\prime})\geq\operatorname{dist}\bigl(Q^{(m)},(Q^{\prime})^{(m)}\bigr)\geq\delta_{m}.$
Also $m\leq n$ and $(\delta_{k})$ is decreasing, so $\delta_{m}\geq\delta_{n}$. Therefore $\operatorname{dist}(Q,Q^{\prime})\geq\delta_{n}$.
Pick $x_{Q}\in Q$ for each $Q\in\mathcal{E}_{n}$. Then $\{x_{Q}\}$ is $\delta_{n}$–separated in $0,1]^{d}$. By Lemma 4.3,
$N_{n}\leq C\,\delta_{n}^{-d}$
for some $C>0$ depending only on $d$. Thus
$r_{1}\cdots r_{n-1}\leq\frac{C}{r_{n}\delta_{n}^{d}}=\frac{C}{L_{n}^{d}}.$
If $r_{k}\geq 2$ for all $k$, then $r_{1}\cdots r_{n-1}\geq 2^{n-1}$, so $L_{n}^{d}\leq C2^{-(n-1)}$ and $L_{n}\to 0$. Hence $\log L_{n}\to\infty$ and, for large $n$,
$\frac{\log(r_{1}\cdots r_{n-1})}{-\log L_{n}}\leq d+\frac{\log C}{-\log L_{n}}.$
Therefore
$\liminf_{n\to\infty}\frac{\log(r_{1}\cdots r_{n-1})}{-\log L_{n}}\leq d.$
Lemma 4.5 (Mass distribution principle).
Let $F\subset\mathbb{R}^{d}$ be a Borel set. Assume there exist $s\geq 0$, constants $C>0$ and $r_{0}>0$, and a Borel probability measure $\mu$ supported on $F$ such that
$\mu\bigl(B(x,r)\bigr)\leq C\,r^{s}\qquad\text{for all }x\in\mathbb{R}^{d},\ 0<r\leq r_{0}.$
Then $\dim_{H}(F)\geq s$.
Proof.
This is the standard mass distribution principle; see [7, Proposition 4.2]. ∎
Now we give Proof of Lemma 4.2:
Proof.
Write $\mathcal{E}_{n}$ for the family of $n$-level cubes, and set $R_{n}:=r_{1}\cdots r_{n}$.
If some parent has more than $r_{n}$ children, choose and fix exactly $r_{n}$ children inside it. Do this for all $n$. This gives a Moran subset $F^{\prime}\subset F$ with the same parameters $(r_{n}),(\delta_{n})$. Then $\dim_{H}F\geq\dim_{H}F^{\prime}$. So we may assume each parent has exactly $r_{n}$ children. Hence $\mathcal{E}_{n}=R_{n}$.
Define a probability measure $\mu$ on $F$ by
$\mu(Q)=R_{n}^{-1}\qquad(Q\in\mathcal{E}_{n}).$
This is consistent by nesting, so $\mu$ is well defined and supported on $F$.
Fix $0<r<\delta_{1}/2$ and set $U:=B(x,r)$. Then $\operatorname{diam}(U)=2r$. Choose $n\geq 1$ such that $\delta_{n}\leq 2r<\delta_{n-1}$. Let $N(U)$ be the number of cubes in $\mathcal{E}_{n}$ that meet $U$.
We first note that $\mathcal{E}_{n-1}$ is $\delta_{n-1}$–separated. Indeed, take distinct $I,I^{\prime}\in\mathcal{E}_{n-1}$ and let $m$ be the first level where their ancestors split. Then their $m$-level ancestors are siblings, so their distance is at least $\delta_{m}\geq\delta_{n-1}$ since $(\delta_{k})$ decreases. Thus $\operatorname{dist}(I,I^{\prime})\geq\delta_{n-1}$.
Hence $U$ meets at most one (n1)(n\!-\!1)-level cube (since $\operatorname{diam}(U)<\delta_{n-1}$). So all $n$-level cubes meeting $U$ lie in one parent, and therefore $N(U)\leq r_{n}$.
Also pick one point $x_{Q}\in Q\cap U$ for each such cube $Q$. Then the set $\{x_{Q}\}$ is $\delta_{n}$–separated in $U$, so by Lemma 4.3,
$N(U)\leq C_{d}\Bigl(\frac{\operatorname{diam}(U)}{\delta_{n}}+1\Bigr)^{d}\leq C\,\Bigl(\frac{r}{\delta_{n}}\Bigr)^{d}$
for a constant $C$ depending only on $d$ (since $\delta_{n}\leq 2r$).
Thus
$N(U)\leq\min\Bigl\{r_{n},\;C\Bigl(\frac{r}{\delta_{n}}\Bigr)^{d}\Bigr\}.$
Fix $s\in(0,d]$ and put $\theta:=s/d\in(0,1]$. Then $\min\{A,B\}\leq A^{1-\theta}B^{\theta}$, so
$N(U)\leq r_{n}^{1-s/d}\cdot\Bigl(C\Bigl(\frac{r}{\delta_{n}}\Bigr)^{d}\Bigr)^{s/d}=C^{s/d}\,r^{s}\,r_{n}^{1-s/d}\,\delta_{n}^{-s}.$
Since each $n$-cube has $\mu$-mass $R_{n}^{-1}$, we get
$\mu(U)\leq R_{n}^{-1}N(U)\leq C^{s/d}\,r^{s}\cdot R_{n}^{-1}r_{n}^{1-s/d}\delta_{n}^{-s}=C^{s/d}\,r^{s}\cdot\frac{1}{R_{n-1}(r_{n}^{1/d}\delta_{n})^{s}}.$
Let
$\alpha:=\liminf_{n\to\infty}\frac{\log R_{n-1}}{-\log(r_{n}^{1/d}\delta_{n})}.$
Take any $s<\alpha$ with $s\leq d$. Then for all large $n$,
$R_{n-1}(r_{n}^{1/d}\delta_{n})^{s}\geq 1.$
For such $r$ (hence such $n$) we have $\mu(B(x,r))\leq C^{\prime}r^{s}$, with $C^{\prime}$ depending only on $d$ and $s$. By Lemma 4.5, $\dim_{H}(F)\geq s$.
Since this holds for all $s<\alpha$ with $s\leq d$, we obtain
$\dim_{H}(F)\geq\min\{d,\alpha\}.$
Finally, Lemma 4.4 gives $\alpha\leq d$, so $\min\{d,\alpha\}=\alpha$, which is exactly (26). ∎
5 Three-stage construction in $\mathbb{N}^{d}$
In this section we construct Moran–type seed sets in $\mathbb{N}^{d}$ and derive lower bounds for their Hausdorff dimension in terms of the polynomial density exponent of the underlying digit set. The construction is intentionally flexible: no divergence or anisotropy assumptions are imposed at this stage, and such conditions will be enforced later only when required by the transference arguments.
The argument proceeds in three steps: we first extract a relatively thin subset $S^{\ast}\subset S$ with controlled growth, then build an extreme seed Moran fractal from $S^{\ast}$, and finally insert digits from a target set along carefully chosen indices while preserving Hausdorff dimension via an almost–Lipschitz elimination map.
5.1 Construction of a relatively thin subset
We define a slowly increasing function $\nu:[1,\infty)\to\mathbb{N}$ by the factorial blocks:
$\nu(\xi)=k\quad\text{for }\xi\in[k!,(k+1)!),\quad k\in\mathbb{N}.$ (27)
Lemma 5.1 ([24], Lemma 3.3).
For every $t>e$, there exists a $\lambda>1$, depending on $t$, such that
$\nu(t^{n})\leq(n\log t)^{\lambda}\quad\text{for all }n\in\mathbb{N}.$
We first show that from any infinite subset of $\mathbb{N}^{d}$ we can extract a subset of zero relative density with a slowly varying growth rate.
Lemma 5.2 (Extraction of a thin subset in $\mathbb{N}^{d}$).
Let $S\subset\mathbb{N}^{d}$ be infinite. Then there exists $S^{\ast}\subset S$ such that:
  • (i)
    $S^{\ast}$ has no “nearest neighbour” pairs in the sense that whenever $\mathbf{v},\mathbf{w}\in S^{\ast}$ with $\mathbf{v}\neq\mathbf{w}$, one has $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$;
  • (ii)
    there exist constants $C_{1},C_{2}>0$ (depending only on $d$) such that for all large $N$,
    $\frac{C_{1}}{\nu(\#(S\cap Q_{N}))}\leq\frac{\#(S^{\ast}\cap Q_{N})}{\#(S\cap Q_{N})}\leq\frac{C_{2}}{\nu(\#(S\cap Q_{N}))};$
  • (iii)
    in particular $\overline{d}(S^{\ast}\mid S)=0$.
Remark 5.3.
The selection is essentially one–dimensional. We first pass to a 22–separated subset $S_{\mathrm{sep}}\subset S$ (in $\|\cdot\|_{\infty}$); this is used later to ensure separation of cylinders in the Moran construction (hence diameter and gap lower bounds). We then enumerate $S_{\mathrm{sep}}$ by nondecreasing $\|\cdot\|_{\infty}$ and apply the same factorial–block thinning rule as in the one–dimensional case. The estimates depend only on the counting function $N_{S_{\mathrm{sep}}}(R)$ and not on the particular enumeration.
Proof.
Let us define the counting function $N_{S}(R)=\#\{\mathbf{v}\in S:\|\mathbf{v}\|_{\infty}\leq R\}$ for $R>0$, and similarly for other subsets.
We first construct a subset $S_{\text{sep}}\subset S$ that satisfies the separation condition. We define a greedy selection process: order the elements of $S$ by their $\ell_{\infty}$ norm: $S=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots\}$.
  • Select $\mathbf{x}_{1}$.
  • Recursively, select the next $\mathbf{x}_{j}$ only if $\|\mathbf{x}_{j}-\mathbf{x}_{i}\|_{\infty}\geq 2$ for all previously selected $\mathbf{x}_{i}$.
Let the resulting set be $S_{\text{sep}}$. Since the number of integer points in an $\ell_{\infty}$-ball of radius 2 is bounded by $K_{d}=3^{d}$, at most $K_{d}$ points from $S$ are discarded for every point selected. Thus, the density of $S_{\text{sep}}$ is comparable to $S$:
$\frac{1}{K_{d}}N_{S}(R)\leq N_{S_{\text{sep}}}(R)\leq N_{S}(R).$ (28)
We now linearize $S_{\text{sep}}$ by enumerating its elements in increasing order of their $\ell_{\infty}$ norm:
$S_{\text{sep}}=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\dots\}.$
Define the index set
$Q=\bigcup_{k=2}^{\infty}\{k!+ik:i=0,1,\dots,k!-1\}$
and construct the final set $S^{\ast}$ by selecting vectors from $S_{\text{sep}}$ with indices in $Q$:
$S^{\ast}=\{\mathbf{u}_{n}\in S_{\text{sep}}:n\in Q\}.$
Verification of properties:
(i) Since $S^{\ast}\subset S_{\text{sep}}$, and $S_{\text{sep}}$ is 2-separated by construction, $S^{\ast}$ satisfies $\|\mathbf{u}-\mathbf{v}\|_{\infty}\geq 2$ for any distinct $\mathbf{u},\mathbf{v}\in S^{\ast}$.
(ii) Let $\xi=N_{S_{\text{sep}}}(R)$ be the number of elements in $S_{\text{sep}}$ up to norm $R$. The number of elements in $S^{\ast}$ up to norm $R$ is exactly $(Q\cap[1,\xi])$. When $\xi\in[k!,(k+1)!)$, we estimate the size of $Q\cap[1,\xi]$ as follows:
Note that by the construction of $Q$, in each interval $[j!,(j+1)!)$ from $j=2$ to j=k1j=k-1, the set $Q$ containins exactly $j!$ elements. We also note, $Q\cap[k!,(k+1)!)$ consists of a partial block containing $\left\lfloor\frac{\xi-k!}{k}\right\rfloor+1$ elements. We decompose $1,\xi]$ into three parts:
$\#(Q\cap[1,\xi])$ $=\#(Q\cap[1,(k-1)!))+\#(Q\cap[(k-1)!,k!))+\#(Q\cap[k!,\xi])$
$\leq\#(Q\cap[1,(k-1)!))+\frac{\#([(k-1)!,k!)\cap\mathbb{N})}{k-1}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}$
$\leq(k-1)!+(k-1)!+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}\leq\frac{3\xi}{\nu(\xi)},$
where we used that $(k-1)!\leq\xi/k$ for sufficiently large $k$ (since $\xi\geq k!$ and $k!(k)=(k+1)!-k!>k\cdot k!$ for $k\geq 2$).
For the lower bound, we have:
$\#(Q\cap[1,\xi])$ $\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{i}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}-1$
$\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{k}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}-1$
$=\frac{\#([2,\xi]\cap\mathbb{N})}{k}-1\geq\frac{\xi}{2\nu(\xi)},$
where the last inequality holds for sufficiently large $\xi$ since $([2,\xi]\cap\mathbb{N})=\lfloor\xi\rfloor-1\geq\xi-2$ and $\frac{\xi-2}{k}-1\geq\frac{\xi}{2k}$ for $\xi\geq 2k+4$, which is satisfied for sufficiently large $\xi$ given that $\xi\geq k!$ and $k!$ grows faster than $2k+4$.
Substituting $k=\nu(\xi)=\nu(N_{S_{\text{sep}}}(R))$ yields:
$\frac{N_{S_{\text{sep}}}(R)}{2\nu(N_{S_{\text{sep}}}(R))}\leq N_{S^{\ast}}(R)\leq\frac{3N_{S_{\text{sep}}}(R)}{\nu(N_{S_{\text{sep}}}(R))}.$
Dividing by $N_{S}(R)$ and using the bounds from (28), we obtain:
$\frac{1}{2K_{d}\cdot\nu(N_{S_{\text{sep}}}(R))}\leq\frac{N_{S^{\ast}}(R)}{N_{S}(R)}\leq\frac{3}{\nu(N_{S_{\text{sep}}}(R))}.$
Since $\nu$ is constant on intervals $[k!,(k+1)!)$ and $N_{S_{\text{sep}}}(R)$ differs from $N_{S}(R)$ by at most a constant factor $K_{d}$, for sufficiently large $R$ we have $\nu(N_{S}(R))-1\leq\nu(N_{S_{\text{sep}}}(R))\leq\nu(N_{S}(R))$. Thus, with $C_{1}=1/(2K_{d})$ and $C_{2}=4$, we obtain the desired bounds.
(iii) The upper bound in (ii) implies that $\frac{\#(S^{\ast}\cap Q_{N})}{\#(S\cap Q_{N})}\to 0$ as $N\to\infty$, hence $\overline{d}(S^{\ast}\mid S)=0$. ∎
This construction ensures all required properties while maintaining the necessary density controls through careful block partitioning and residue class selection.
5.2 Extreme seed sets and their dimension
We begin by isolating the ambient “no–repetition” sets in which our seed (root) constructions will take place. For $S\subset\mathbb{N}^{d}$ define
$E_{S}^{\mathrm{vec}}:=\Bigl\{\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}:\mathbf{a}_{n}(\mathbf{x})\in S\ \text{for all }n\geq 1,\ \text{and }\mathbf{a}_{m}(\mathbf{x})\neq\mathbf{a}_{n}(\mathbf{x})\ \text{for all }m\neq n\Bigr\},$
where $\mathbf{a}_{n}(\mathbf{x}):=(a_{n}(x_{1}),\dots,a_{n}(x_{d}))\in\mathbb{N}^{d}$ denotes the nn–th digit vector of $\mathbf{x}$.
For parameters $t>L>1$ with $Lt<t^{2}$, any $\mathbf{x}\in R_{t,L}(K)$ satisfies
$t^{n}\leq\|\mathbf{a}_{n}(\mathbf{x})\|_{\infty}<Lt^{n}\qquad(n\geq 1).$
Since the intervals $t^{n},Lt^{n})$ are pairwise disjoint (equivalently $Lt^{n}<t^{n+1}$), we have $\mathbf{a}_{m}(\mathbf{x})\neq\mathbf{a}_{n}(\mathbf{x})$ whenever $m\neq n$. In particular, if $K\subset S$, then
$R_{t,L}(K)\subset E_{S}^{\mathrm{vec}}.$
We also use the divergence subspace
$\mathcal{E}_{\mathbb{N}^{d}}:=\Bigl\{\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}:a_{n}(x_{j})\to\infty\ \text{as }n\to\infty\ \text{for each }1\leq j\leq d\Bigr\},$
and recall the digit–restricted divergence set
$\mathcal{E}_{S}^{\mathrm{vec}}=E_{S}^{\mathrm{vec}}\cap\mathcal{E}_{\mathbb{N}^{d}}.$
Finally, if $S\subset\mathcal{C}_{K}$ is uniformly KK–balanced, then for any $\mathbf{a}_{n}(\mathbf{x})\in S$ we have $\min_{j}a_{n}(x_{j})\geq K^{-1}\|\mathbf{a}_{n}(\mathbf{x})\|_{\infty}$. Hence $\|\mathbf{a}_{n}(\mathbf{x})\|_{\infty}\to\infty$ implies $\mathbf{x}\in\mathcal{E}_{\mathbb{N}^{d}}$, so in the balanced setting we may regard $\mathcal{E}_{S}^{\mathrm{vec}}$ as the natural ambient space for the seed constructions.
Remark 5.4.
In what follows we construct seed sets inside $E_{d}^{\mathrm{vec}}$. Later, when restricting digits to a subset $S\subset\mathcal{C}_{K}$, we may further intersect with $\mathcal{E}_{\mathbb{N}^{d}}$ (or with $\mathcal{E}_{S}$) in order to enforce divergence of digits. We do not impose additional coordinatewise restrictions beyond what is required. In [24], the fractal transference principle is formulated on $\mathcal{E}$ for $d=1$; for consistency with that framework, we will also work on $\mathcal{E}_{\mathbb{N}^{d}}$ in the final arguments.
Given the set $S^{\ast}$ obtained in Lemma 5.2, we now construct Moran-type subsets of $E_{d}^{\mathrm{vec}}$ whose digit vectors lie in $S^{\ast}$ and grow along a prescribed scale. These sets will serve as extreme seed sets, achieving the maximal Hausdorff dimension permitted by the growth properties of $S$.
Definition 5.5 (Seed sets associated with $S$).
Fix parameters $t>L>1$, with $t$ sufficiently large. For an infinite subset $K\subset S$ we define
$R_{t,L}(K):=\Bigl\{\mathbf{x}\in E_{d}^{\mathrm{vec}}:\mathbf{a}_{n}(\mathbf{x})\in K\cap\bigl\{\mathbf{v}\in\mathbb{N}^{d}:\ t^{n}\leq\|\mathbf{v}\|_{\infty}<Lt^{n}\bigr\}\ \text{for all }n\geq 1\Bigr\}.$
Non-empty sets of this form are called seed sets.
Definition 5.6 (Extreme seed sets).
Let $S\subset\mathbb{N}^{d}$ be infinite and let $R_{t,L}(K)$ be a seed set constructed from some $K\subset S$. We say that $R_{t,L}(K)$ is an extreme seed set associated with $S$ if the following conditions hold:
  • (A1)
    $S$ has polynomial density of exponent $\alpha\geq 1$;
  • (A2)
    $K$ contains no nearest neighbour pairs;
  • (A3)
    $\overline{d}(K\mid S)=0$;
  • (A4)
    $\dim_{H}R_{t,L}(K)\geq d/(2\alpha)$.
The next proposition is a multidimensional adaptation of the extreme–seed construction in the one–dimensional setting of [24].
Proposition 5.7 (Existence of extreme seed sets).
Let $S\subset\mathbb{N}^{d}$ have polynomial density of exponent $\alpha\geq 1$. Then there exist a subset $S^{\ast}\subset S$ and parameters $t>L>1$, with $t$ sufficiently large, such that $R_{t,L}(S^{\ast})$ is an extreme seed set associated with $S$. Moreover, $S^{\ast}$ can be chosen relatively sparse in $S$, in the sense that
$\overline{d}(S^{\ast}\mid S)=0.$ (29)
If in addition $S\subset\mathcal{C}_{K}$ for some $K\geq 1$, then the lower bound in (A4) is sharp and in fact
$\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2\alpha}.$ (30)
In this balanced case, one may also redefine the seed set by restricting the ambient space to $\mathcal{E}_{\mathbb{N}^{d}}$, that is, by replacing $E_{d}^{\mathrm{vec}}$ with $\mathcal{E}_{\mathbb{N}^{d}}$ in the definition of $R_{t,L}(K)$, without affecting any of the above conclusions.
Remark 5.8.
Proposition 5.7 is a higher–dimensional adaptation of the one–dimensional extreme–seed construction in [24]. Polynomial density alone does not force coordinatewise divergence of digits: for example, $S=\mathbb{N}\times F$ with $F\subset\mathbb{N}^{d-1}$ finite has exponent $\alpha=1$ but yields non–divergent seed sets.
In the later transference arguments we impose divergence only when it is needed, either by working with (upper) density / upper Banach density witnesses whose translations drift to infinity, or by assuming a uniform balance condition $S\subset\mathcal{C}_{K}$. In the balanced case, the dimension lower bound in Proposition 5.7 is sharp, giving (30). We also note that stronger growth assumptions (e.g. $|S\cap Q_{N}|\gg N^{d-1+\varepsilon}$) would already force divergence, but we do not use this here.
Proof of Proposition 5.7.
We work with the notion of extreme seed sets introduced in Definition 5.6.
Assume that there exist $S^{\ast}\subset S$ and parameters $t>L>1$ such that $R_{t,L}(S^{\ast})$ is an extreme seed set associated with $S$, that is, (A2)–(A4) hold with $K=S^{\ast}$ and $\overline{d}(S^{\ast}\mid S)=0$. Condition (A1) is part of the hypothesis.
If in addition $S\subset\mathcal{C}_{K}$, then Corollary 3.5 yields the upper bound
$\dim_{H}(\mathcal{E}_{S})\leq\frac{d}{2\alpha}.$
Since $R_{t,L}(S^{\ast})\subset\mathcal{E}_{S}$ by construction and (A4) provides the matching lower bound, we obtain
$\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2\alpha},$
which proves the “Moreover” statement of the proposition.
In the balanced case, the uniform balance condition implies that the minimal digit in each coordinate diverges along the construction of $R_{t,L}(S^{\ast})$. Consequently,
$R_{t,L}(S^{\ast})\subset\mathcal{E}_{\mathbb{N}^{d}},$
and the seed set may equivalently be defined with ambient space $\mathcal{E}_{\mathbb{N}^{d}}$ in place of $E_{d}^{\mathrm{vec}}$, without affecting any of the conclusions.
It therefore remains to establish the existence of a subset $S^{\ast}\subset S$ and parameters $t>L>1$ (with $t$ sufficiently large) such that $R_{t,L}(S^{\ast})$ satisfies (A2)–(A4) and $\overline{d}(S^{\ast}\mid S)=0$.
Step 1: Extraction of a thin subset $S^{\ast}$.
Since $S\subset\mathbb{N}^{d}$ has polynomial density with exponent $\alpha\geq 1$, there exist constants $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $N$,
$C_{1}\frac{N^{d/\alpha}}{(\log N)^{\beta}}\leq|S\cap Q_{N}|\leq C_{2}\frac{N^{d/\alpha}}{(\log N)^{\beta}}.$
By Lemma 5.2, there exists $S^{\ast}\subset S$ such that:
  • (i)
    $S^{\ast}$ has no nearest neighbour pairs: $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$ for all distinct $\mathbf{v},\mathbf{w}\in S^{\ast}$;
  • (ii)
    There exist constants $c_{1},c_{2}>0$ such that for all large $N$,
    $\frac{c_{1}}{\nu(|S\cap Q_{N}|)}\leq\frac{|S^{\ast}\cap Q_{N}|}{|S\cap Q_{N}|}\leq\frac{c_{2}}{\nu(|S\cap Q_{N}|)};$
  • (iii)
    In particular, $\overline{d}(S^{\ast}\mid S)=0$.
Thus (A2) and (A3) hold for $K=S^{\ast}$.
Step 2: Choice of parameters and growth estimates.
By Lemma 2.10, for any $t>L>1$ there exist constants $C^{\prime}_{1},C^{\prime}_{2}>0$, independent of $t$ and $L$, such that for all sufficiently large $n$,
$C^{\prime}_{1}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq C^{\prime}_{2}L^{-d/\alpha}.$
Choose $L>1$ so that
$\frac{c_{1}C_{2}^{\prime-1}L^{d/\alpha}}{2}-c_{2}>C_{1}^{-1},$
and then fix $t>\max\{L,2\}$ sufficiently large so that $\nu(|S\cap Q_{t}|)\geq 2$.
For each $n\in\mathbb{N}$, define the annulus
$A_{n}:=\{\mathbf{v}\in S^{\ast}:\ t^{n}\leq\|\mathbf{v}\|_{\infty}<Lt^{n}\},\qquad r_{n}:=|A_{n}|.$
Fix $n$ sufficiently large and choose $k=k(n)$ such that
$k!\leq|S\cap Q_{t^{n}}|<(k+1)!.$
By the left-hand inequality above, for all sufficiently large $n$,
$|S\cap Q_{Lt^{n}}|\leq(C^{\prime}_{1})^{-1}L^{d/\alpha}\,|S\cap Q_{t^{n}}|<(k+2)!,$
after increasing $n$ if necessary.
By the definition of $\nu$, this implies:
$0\leq\nu(|S\cap Q_{Lt^{n}}|)-\nu(|S\cap Q_{t^{n}}|)\leq 1.$
By Lemma 5.1, there exists $\lambda>1$ such that
$\nu(|S\cap Q_{t^{n}}|)\leq\nu(t^{nd})\leq(nd\log t)^{\lambda}=d^{\lambda}(n\log t)^{\lambda}$
for all $n\geq 1$.
Since we can choose $\lambda^{\prime}>\lambda$ such that $d^{\lambda}(n\log t)^{\lambda}\leq(n\log t)^{\lambda^{\prime}}$ for all $n$, we may therefore assume without loss of generality that
$\nu(|S\cap Q_{t^{n}}|)\leq(n\log t)^{\lambda}\quad\text{for all }n.$
We now estimate $r_{n}$:
$r_{n}$ $=|S^{\ast}\cap Q_{Lt^{n}}|-|S^{\ast}\cap Q_{t^{n}}|$
$\geq\frac{c_{1}|S\cap Q_{Lt^{n}}|}{\nu(|S\cap Q_{Lt^{n}}|)}-\frac{c_{2}|S\cap Q_{t^{n}}|}{\nu(|S\cap Q_{t^{n}}|)}$
$\geq\frac{c_{1}{C^{\prime}_{2}}^{-1}L^{d/\alpha}|S\cap Q_{t^{n}}|}{\nu(|S\cap Q_{t^{n}}|)+1}-\frac{c_{2}|S\cap Q_{t^{n}}|}{\nu(|S\cap Q_{t^{n}}|)}$
$=\frac{|S\cap Q_{t^{n}}|}{\nu(|S\cap Q_{t^{n}}|)}\left(\frac{c_{1}{C^{\prime}_{2}}^{-1}L^{d/\alpha}}{1+\frac{1}{\nu(|S\cap Q_{t^{n}}|)}}-c_{2}\right)$
$\geq\frac{|S\cap Q_{t^{n}}|}{\nu(|S\cap Q_{t^{n}}|)}\left(\frac{c_{1}{C^{\prime}_{2}}^{-1}L^{d/\alpha}}{2}-c_{2}\right).$
where the last inequality follows from $\nu(|S\cap Q_{t^{n}}|)\geq 2$ for large $n$.
By our choice of $L$, we have $\frac{c_{1}{C^{\prime}_{2}}^{-1}L^{d/\alpha}}{2}-c_{2}\geq{C_{1}}^{-1}$, so
$r_{n}\geq\frac{|S\cap Q_{t^{n}}|}{{C_{1}}\nu(|S\cap Q_{t^{n}}|)}.$
Step 3: Construction of the Moran fractal.
Using the bounds $|S\cap Q_{t^{n}}|\geq C_{1}t^{nd/\alpha}/(n\log t)^{\beta}$ and $\nu(|S\cap Q_{t^{n}}|)\leq(n\log t)^{\lambda}$, we obtain
$r_{n}\geq\frac{t^{nd/\alpha}}{(n\log t)^{\beta}}\cdot\frac{1}{(n\log t)^{\lambda}}=\frac{t^{nd/\alpha}}{(n\log t)^{\rho}},$ (31)
where $\rho=\beta+\lambda$.
Consider the symbolic space:
$\Sigma^{\ast}=\{(\mathbf{v}_{1},\mathbf{v}_{2},\ldots):\mathbf{v}_{n}\in A_{n}\text{ for all }n\}$
and
$\Sigma^{\ast}_{n}=\{(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}):\mathbf{v}_{i}\in A_{i}\text{ for all }i\in\{1,\ldots,n\}\}.$
For each $\omega=(\mathbf{v}_{1},\ldots,\mathbf{v}_{n})\in\Sigma^{\ast}_{n}$, define the cylinder set:
$I_{\omega}=\{\mathbf{x}\in(0,1)^{d}:\mathbf{a}_{i}(\mathbf{x})=\mathbf{v}_{i}\text{ for }i=1,\ldots,n\}.$
Define
$F:=\bigcap_{n=1}^{\infty}\bigcup_{\omega\in\Sigma_{n}^{\ast}}\overline{I_{\omega}}.$
We claim that $F\subset R_{t,L}(S^{\ast})$. Fix $\mathbf{x}\in F$ and choose for each $n$ a word $\omega\in\Sigma_{n}^{\ast}$ such that $\mathbf{x}\in\overline{I_{\omega}}$. If some coordinate $x_{j}\in\mathbb{Q}$, then $x_{j}$ would have to be an endpoint of a one–dimensional cylinder of arbitrarily large order. However, cylinder endpoints are convergents $p_{n}/q_{n}$, and the corresponding denominators satisfy $q_{n}\to\infty$ (as $q_{n+1}=a_{n+1}q_{n}+q_{n-1}\geq q_{n}+q_{n-1}$). Hence a fixed rational number can be an endpoint for only finitely many orders, a contradiction. Thus $\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}$.
For such $\mathbf{x}$, membership in $\overline{I_{\omega}}$ implies $\mathbf{x}\in I_{\omega}$, and the length-$n$ word $\omega\in\Sigma_{n}^{\ast}$ with $\mathbf{x}\in I_{\omega}$ is unique. Therefore these words are consistent in $n$ and define an infinite admissible digit sequence in $S^{\ast}$, so $\mathbf{x}\in R_{t,L}(S^{\ast})$.
We next verify that $F$ is a Moran fractal. For each $n\geq 1$ set $\mathcal{Q}_{n}:=\{\overline{I_{\omega}}:\omega\in\Sigma_{n}^{\ast}\}$. By Remark 2.9,
$\max_{Q\in\mathcal{Q}_{n}}\operatorname{diam}(Q)=\max_{\omega\in\Sigma_{n}^{\ast}}\operatorname{diam}(\overline{I_{\omega}})\leq\sqrt{d}\,C\,\vartheta^{n}\to 0,$
so the vanishing mesh condition holds.
Since $t>L$, the annuli $\{v\in\mathbb{N}^{d}:t^{n}\leq\|v\|_{\infty}<Lt^{n}\}$ are pairwise disjoint. Fix $n\geq 2$ and a parent cylinder $\overline{I_{\omega}}$ with $\omega\in\Sigma_{n-1}^{\ast}$. By (31), the number of children $\overline{I_{\omega^{\prime}}}\subset\overline{I_{\omega}}$ at level $n$ is at least $\lfloor t^{nd/\alpha}/(n\log t)^{\rho}\rfloor$.
To obtain separation inside a parent, let $\omega,\omega^{\prime}\in\Sigma_{n}^{\ast}$ be distinct words with the same prefix of length n1n-1. Since distinct vectors in $S^{\ast}$ satisfy $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$, Lemma 2.7 yields
$\operatorname{dist}(I_{\omega},I_{\omega^{\prime}})\geq c\,\min_{1\leq j\leq d}\min\bigl(|I_{\omega}^{(j)}|,\ |I_{\omega^{\prime}}^{(j)}|\bigr).$
For any $\omega\in\Sigma_{n}^{\ast}$, Lemma 2.1(i) gives
$|I_{\omega}^{(j)}|\geq\frac{1}{2}\prod_{i=1}^{n}(a_{i}^{(j)}+1)^{-2}.$
Using $a_{i}^{(j)}\leq\|\mathbf{v}_{i}\|_{\infty}\leq Lt^{i}$ and $Lt^{i}+1\leq 2Lt^{i}$, we obtain
$|I_{\omega}^{(j)}|\geq 2^{-2n-1}L^{-2n}t^{-n(n+1)}.$
Hence, for a constant $c^{\prime}>0$ depending only on $d$,
$\operatorname{dist}(I_{\omega},I_{\omega^{\prime}})\geq\delta_{n},\qquad\delta_{n}:=c^{\prime}(2L)^{-2n}t^{-n(n+1)}.$
Therefore $F$ is a Moran fractal with branching numbers $r_{n}=\lfloor t^{nd/\alpha}/(n\log t)^{\rho}\rfloor$ and gap scales $(\delta_{n})_{n\geq 1}$.
Step 4: Hausdorff dimension estimate.
By Lemma 4.2,
$\dim_{H}F\geq\liminf_{n\to\infty}\frac{\log\bigl(\prod_{i=1}^{n-1}r_{i}\bigr)}{-\log\bigl(r_{n}^{1/d}\delta_{n}\bigr)}.$
Since $\log r_{i}=\frac{d}{\alpha}i\log t+O(\log i)$, we have
$\log\Bigl(\prod_{i=1}^{n-1}r_{i}\Bigr)=\sum_{i=1}^{n-1}\log r_{i}=\frac{d}{2\alpha}n^{2}\log t+o(n^{2}\log t).$
Moreover,
$\log\bigl(r_{n}^{1/d}\delta_{n}\bigr)=-\frac{1}{d}\log r_{n}-\log\delta_{n}=n(n+1)\log t+o(n^{2}\log t)=n^{2}\log t+o(n^{2}\log t),$
because $\log\delta_{n}=n(n+1)\log t+O(n)$ and $\log r_{n}=O(n)$. Therefore,
$\dim_{H}F\geq\frac{d}{2\alpha}.$
Since $F\subset R_{t,L}(S^{\ast})$, it follows that
$\dim_{H}R_{t,L}(S^{\ast})\geq\frac{d}{2\alpha},$
which establishes (A4).
This establishes the existence statement, and together with the discussion at the beginning of the proof, completes the proof of Proposition 5.7.
5.3 Adding digits and elimination maps
Starting from an extreme seed set $R_{t,L}(B)$, we now insert finitely many digit vectors from a target set $A$ along a sparse sequence of levels, so that the resulting digit set realizes the desired (relative) density while the Hausdorff dimension is essentially unchanged. The key technical point is that the associated elimination map, which deletes the inserted blocks, is Hölder with every exponent $\gamma\in(0,1)$ once suitable growth conditions are enforced.
5.3.1 Rigorous digit insertion in higher dimensions
Fix parameters $t>L>1$. Let $A,B\subset\mathbb{N}^{d}$ be infinite sets such that $R_{t,L}(B)\neq\emptyset$. Let $(M_{k})_{k\geq 1}$ be a strictly increasing sequence with $M_{0}=0$, and let $(W_{k})_{k\geq 1}$ be finite subsets of $A$. For each $k$, fix an ordering
$W_{k}=\{w_{k,1},\dots,w_{k,m_{k}}\},\qquad m_{k}:=|W_{k}|.$
Given $\mathbf{y}\in R_{t,L}(B)$ with digit sequence
$\bigl(\mathbf{a}_{1}(\mathbf{y}),\mathbf{a}_{2}(\mathbf{y}),\dots\bigr)\in(\mathbb{N}^{d})^{\mathbb{N}},$
we define $\mathbf{x}=\iota(\mathbf{y})$ by inserting the word $(w_{k,1},\dots,w_{k,m_{k}})$ immediately after $\mathbf{a}_{M_{k}}(\mathbf{y})$ for each $k$, i.e.
$(\mathbf{a}_{n}(\mathbf{x}))_{n\geq 1}=\mathbf{a}_{1}(\mathbf{y}),\dots,\mathbf{a}_{M_{1}}(\mathbf{y}),w_{1,1},\dots,w_{1,m_{1}},\mathbf{a}_{M_{1}+1}(\mathbf{y}),\dots,\mathbf{a}_{M_{2}}(\mathbf{y}),w_{2,1},\dots,w_{2,m_{2}},\ \dots\ .$
We set
$R_{t,L}(B,A,(M_{k}),(W_{k})):=\{\iota(\mathbf{y}):\mathbf{y}\in R_{t,L}(B)\}.$
The elimination map $f$ is defined on this set by deleting the inserted blocks and recovering the underlying seed sequence:
$f(\iota(\mathbf{y})):=\mathbf{y}.$
Thus $f$ is a bijection from $R_{t,L}(B,A,(M_{k}),(W_{k}))$ onto $R_{t,L}(B)$.
Remark 5.9.
In our applications we choose the inserted blocks so that no digit vector is repeated (e.g. by ensuring $B\cap\bigcup_{k}W_{k}=\emptyset$ and that the $W_{k}$ are pairwise disjoint). This guarantees that the inserted set remains inside the natural no–repetition ambient space. The ordering of each $W_{k}$ is irrelevant for all dimension estimates and only serves to make $\iota$ (hence $f$) well-defined.
5.3.2 Growth conditions
Let $(\varepsilon_{k})_{k\geq 1}$ be a decreasing sequence of positive numbers with $\varepsilon_{k}\downarrow 0$. We record the quantitative constraints we will impose (inductively) on $M_{k}$ (and later on $W_{k}$). These conditions are chosen to control how much the insertion can distort continued-fraction cylinders; compare [24, Section 3.4].
$Lt^{M_{k-1}}$ $<M_{k}^{1/(2d)}$ (32)
$Lt^{M_{k-1}}+M_{k}^{1/(2d)}$ $<t^{M_{k}}$ (33)
$\frac{1}{(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1)^{2^{d+1}M_{k}^{1/2}}}$ $\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}t^{-2i}\right)^{3\varepsilon_{k}}$ (34)
$\inf_{n\geq M_{k}}\frac{\prod_{i=1}^{n+1}(t^{i+2})^{-2}\cdot\left(\prod_{i=1}^{n}t^{-2i}\right)^{3\varepsilon_{k}}}{\left(\prod_{i=1}^{n}t^{-2i}\right)^{1+4\varepsilon_{k}}}$ $\geq 2.$ (35)
If the blocks $(W_{k})$ have already been specified, we also require
$\prod_{w\in W_{k}}(\|w\|_{\infty}+1)^{-2}\;\geq\;\left(\prod_{i=M_{k-1}+1}^{M_{k}}t^{-2i}\right)^{3\varepsilon_{k}}\qquad\text{whenever }W_{k}\neq\emptyset.$ (36)
Remark 5.10.
Heuristically, (32)–(33) ensure that the inserted digits are small compared with the seed scale $t^{M_{k}}$, while (34)–(36) quantify that the cumulative contraction contributed by the inserted blocks is dominated by the contraction of the seed levels. This is what ultimately yields the almost–Lipschitz (Hölder for every $\gamma<1$) behaviour of the elimination map.
Lemma 5.11.
Fix $t>L>1$ and a sequence $\varepsilon_{k}\downarrow 0$. Given $M_{k-1}$, one can choose $M_{k}$ sufficiently large so that (32)–(35) hold. In particular, there exists a strictly increasing sequence $(M_{k})_{k\geq 1}$ satisfying (32)–(35).
Proof.
Conditions (32)–(33) are immediate. We verify (34) and (35).
For (34), the right-hand side equals
$t^{-3\varepsilon_{k}(M_{k}+M_{k-1}+1)(M_{k}-M_{k-1})}.$
The left-hand side is
$t^{-\,2^{d+1}M_{k}^{1/2}\,\log_{t}(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1)}.$
Since
$\log_{t}(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1)<M_{k}^{1/2}$
for large $M_{k}$, the left-hand side is $t^{-2^{d+1}M_{k}}$. Thus it suffices to check
$t^{-2^{d+1}M_{k}}\geq t^{-3\varepsilon_{k}(M_{k}+M_{k-1}+1)(M_{k}-M_{k-1})},$
which holds for large $M_{k}$.
For (35), the expression simplifies to
$t^{(n+1)(n\varepsilon_{k}-4)}.$
This exceeds 22 once $n\varepsilon_{k}>4$, which is the case for sufficiently large $M_{k}$. ∎
5.3.3 Almost Lipschitz property with explicit estimates
Theorem 5.12.
Fix parameters $t>L>1$. Let $A,B\subset\mathbb{N}^{d}$ be infinite sets such that $R_{t,L}(B)$ is non-empty and $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$ for all distinct $\mathbf{v},\mathbf{w}\in B$. Let $(M_{k})_{k\geq 1}$ be a strictly increasing sequence with $M_{0}=0$, $(W_{k})_{k\geq 1}$ finite subsets of $A$, and $(\varepsilon_{k})_{k\geq 1}$ a decreasing sequence of positive numbers converging to 0, all satisfying (35) and (36).
Then, for every $\gamma\in(0,1)$ there exists $C_{\gamma}>0$ such that
$\|f(\mathbf{x}_{1})-f(\mathbf{x}_{2})\|_{\infty}\;\leq\;C_{\gamma}\,\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}^{\gamma}\quad\text{for all }\mathbf{x}_{1},\mathbf{x}_{2}\in R_{t,L}(B,A,(M_{k}),(W_{k})).$ (37)
Thus $f$ is Hölder continuous with every exponent $\gamma\in(0,1)$ on this set.
Remark 5.13.
The hypotheses (35)–(36) are the natural higher–dimensional counterparts of the one–dimensional assumptions in [24, Proposition 3.5]. The proof follows the same strategy, with the only changes coming from the geometry of dd–dimensional cylinders (diameter and separation estimates).
Proof of Theorem 5.12.
For $\mathbf{x}_{1},\mathbf{x}_{2}\in R_{t,L}(B,A,(M_{k}),(W_{k}))$, set $\mathbf{y}_{i}=f(\mathbf{x}_{i})\in R_{t,L}(B)$ and let $a_{n}(\mathbf{y}_{i})$ denote their $n$-th digits. Define
$n:=s(\mathbf{y}_{1},\mathbf{y}_{2}):=\min\{m\geq 0:a_{m+1}(\mathbf{y}_{1})\neq a_{m+1}(\mathbf{y}_{2})\}.$
We distinguish whether the first discrepancy occurs before or after the first insertion scale $M_{k}$.
Case 1: $n<M_{k}$. Here the difference occurs before any insertions. Both points lie among finitely many possible pairs of cylinders, and by Proposition 2.5 and Lemma 2.6 there exists $\delta_{0}>0$ such that $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\delta_{0}$. Since $\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq 1$, choosing $C_{\gamma}^{(1)}:=\delta_{0}^{-\gamma}$ gives
$\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq C_{\gamma}^{(1)}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}^{\gamma}.$ (38)
Case 2: $n\geq M_{k}$. Let $q\geq k$ be such that $M_{q}\leq n<M_{q+1}$. By construction,
$\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})\quad\text{for all }1\leq i\leq n+\sum_{j=1}^{q}|W_{j}|.$
Thus the first discrepancy for $(\mathbf{x}_{1},\mathbf{x}_{2})$ occurs at
$n^{\prime}:=n+\sum_{j=1}^{q}|W_{j}|+1,$
with $\mathbf{a}_{n^{\prime}}(\mathbf{x}_{1})=\mathbf{a}_{n+1}(\mathbf{y}_{1})\neq\mathbf{a}_{n+1}(\mathbf{y}_{2})=\mathbf{a}_{n^{\prime}}(\mathbf{x}_{2})$.
Lemma 2.6 gives
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{1}\,\operatorname{diam}(I(\mathbf{a}_{1}(\mathbf{x}_{1}),\dots,\mathbf{a}_{n^{\prime}}(\mathbf{x}_{1})))\cdot\max_{j\in\mathcal{D}}\frac{|a_{n+1}^{(j)}(\mathbf{y}_{2})-a_{n+1}^{(j)}(\mathbf{y}_{1})|}{a_{n+1}^{(j)}(\mathbf{y}_{1})a_{n+1}^{(j)}(\mathbf{y}_{2})},$
where $\mathcal{D}=\{1\leq j\leq d:a_{n+1}^{(j)}(\mathbf{x}_{1})\neq a_{n+1}^{(j)}(\mathbf{x}_{2})\}$. Since $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$ for distinct $\mathbf{v},\mathbf{w}\in B$ and $a_{n+1}^{(j)}(\mathbf{y}_{i})\leq Lt^{n+1}$,
$\max_{j\in\mathcal{D}}\frac{|a_{n+1}^{(j)}(\mathbf{y}_{2})-a_{n+1}^{(j)}(\mathbf{y}_{1})|}{a_{n+1}^{(j)}(\mathbf{y}_{1})a_{n+1}^{(j)}(\mathbf{y}_{2})}\geq(Lt^{n+1})^{-2}.$
Using the lower diameter bound of Proposition 2.5, we have
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{2}(Lt^{n+1})^{-2}\prod_{i=1}^{n}\bigl(\|a_{i}(\mathbf{y}_{1})\|_{\infty}+1\bigr)^{-2}\prod_{j=1}^{q}\prod_{w\in W_{j}}\bigl(\|w\|_{\infty}+1\bigr)^{-2}.$
Since $t$ is chosen sufficiently large, we may assume $t^{2}\geq L+1$, hence $Lt^{i}+1\leq(L+1)t^{i}\leq t^{i+2}$ for all $i\geq 1$. Therefore,
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{2}\prod_{i=1}^{n+1}t^{-2(i+2)}\prod_{j=1}^{q}\prod_{w\in W_{j}}(\|w\|_{\infty}+1)^{-2}.$
By (36), for each $1\leq j\leq q$ we have
$\prod_{w\in W_{j}}(\|w\|_{\infty}+1)^{-2}\geq\left(\prod_{s=M_{j-1}+1}^{M_{j}}t^{-2s}\right)^{3\varepsilon_{j}}.$
Hence
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{2}\prod_{i=1}^{n+1}t^{-2(i+2)}\prod_{j=1}^{q}\left(\prod_{s=M_{j-1}+1}^{M_{j}}t^{-2s}\right)^{3\varepsilon_{j}}.$
Because $(\varepsilon_{j})$ is decreasing and $q\geq k$, we obtain
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{2}\prod_{i=1}^{n+1}t^{-2(i+2)}\left(\prod_{j=1}^{k}\left(\prod_{i=M_{j-1}+1}^{M_{j}}t^{-2i}\right)^{3\varepsilon_{j}}\right)\left(\prod_{i=1}^{M_{q}}t^{-2i}\right)^{3\varepsilon_{k}}.$
Let
$D_{k}:=\prod_{j=1}^{k}\left(\prod_{i=M_{j-1}+1}^{M_{j}}t^{-2i}\right)^{3\varepsilon_{j}}.$
Since $M_{q}\leq n$, we further have
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{2}D_{k}\prod_{i=1}^{n+1}t^{-2(i+2)}\left(\prod_{s=1}^{n}t^{-2s}\right)^{3\varepsilon_{k}}.$
Using (35),
$\inf_{n\geq M_{k}}\frac{\prod_{i=1}^{n+1}t^{-2(i+2)}\cdot\left(\prod_{i=1}^{n}t^{-2i}\right)^{3\varepsilon_{k}}}{\left(\prod_{i=1}^{n}t^{-2i}\right)^{1+4\varepsilon_{k}}}\geq 2,$
we conclude
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq 2C_{2}D_{k}\,\left(\prod_{i=1}^{n}t^{-2i}\right)^{1+4\varepsilon_{k}}.$
On the other hand, the upper cylinder diameter bound gives
$\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq C_{3}\prod_{i=1}^{n}t^{-2i}.$
Combining,
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\frac{2C_{2}D_{k}}{C_{3}^{1/(1+4\varepsilon_{q})}}\,\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}^{1+4\varepsilon_{q}},$
which implies
$\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq C_{*}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}^{\theta_{q}},\qquad\theta_{q}:=\frac{1}{1+4\varepsilon_{q}}.$ (39)
Fix $\gamma\in(0,1)$. Since $\varepsilon_{k}\to 0$, there exists $K$ with $\theta_{k}\geq\gamma$ for all $k\geq K$. If $n\geq M_{K}$, then $q\geq K$ and (39) gives
$\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq C_{4}\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}^{\gamma}.$
If $n<M_{K}$, applying (38) (with $M_{1}$ replaced by $M_{K}$) yields the same inequality. With
$C_{\gamma}:=\max\{C_{\gamma}^{(1)},C_{4}\},$
we obtain
$\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq C_{\gamma}\,\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}^{\gamma}.$
This completes the proof. ∎
Corollary 5.14.
With the above construction:
$\dim_{H}R_{t,L}(B,A,(M_{k}),(W_{k}))\geq\dim_{H}R_{t,L}(B)$
Proof.
This follows immediately from Theorem 5.12 and Lemma 2.15. ∎
6 Proofs of the main theorems
6.1 Proof of Theorem 3.1
We now give the proof of Theorem 3.1, which relies on Proposition 5.7 to establish the upper bound.
Proof of Theorem 3.1.
Upper bound. Fix $\varepsilon>0$ and choose $N\geq N_{0}(\varepsilon)$ as in Theorem 3.3 such that
$\dim_{H}(\widetilde{\mathcal{E}}_{N})\leq\frac{d}{2}+\varepsilon,$
where
$E_{\geq N}:=\{x\in(0,1):a_{n}(x)\geq N\text{ for all }n\geq 1\},\qquad\widetilde{\mathcal{E}}_{N}:=(E_{\geq N})^{d}\subset(0,1)^{d}.$
Let $\mathbf{x}\in\mathcal{E}_{S}$. Since $a_{n}(x_{j})\to\infty$ for each $1\leq j\leq d$, there exists $n_{0}$ such that $a_{n}(x_{j})\geq N$ for all $n\geq n_{0}$ and all $j$. Equivalently, $\mathbf{x}$ belongs to the set obtained from $\widetilde{\mathcal{E}}_{N}$ by modifying only finitely many initial digit vectors. By Lemma 3.7,
$\dim_{H}(\mathcal{E}_{S})\leq\dim_{H}(\widetilde{\mathcal{E}}_{N})\leq\frac{d}{2}+\varepsilon.$
The same argument applies to $\mathcal{E}_{S}^{\mathrm{vec}}$ (in particular, $\mathcal{E}_{S}^{\mathrm{vec}}\subset\mathcal{E}_{S}$), and hence
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\leq\frac{d}{2}+\varepsilon.$
Letting $\varepsilon\downarrow 0$ yields
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\leq\dim_{H}(\mathcal{E}_{S})\leq\frac{d}{2}.$
Lower bound. Assume $\overline{d}(S)>0$. By Lemma 2.13 there exists $K\geq 1$ such that
$\overline{d}\bigl(S\cap\mathcal{C}_{K}\bigr)>0,\qquad\mathcal{C}_{K}=\Bigl\{\mathbf{a}\in\mathbb{N}^{d}:\frac{\max_{j}a_{j}}{\min_{j}a_{j}}\leq K\Bigr\}.$
Set $S_{K}:=S\cap\mathcal{C}_{K}$. Then $S_{K}$ is infinite and uniformly KK–balanced. Moreover, for every $N\geq 1$ we have $|S_{K}\cap Q_{N}|\leq|Q_{N}|=N^{d}$, so $S_{K}$ has polynomial density with exponent $\alpha=1$ in the sense of Proposition 5.7.
We may therefore apply Proposition 5.7 to $S_{K}$ with $\alpha=1$. It yields a subset $S^{\ast}\subset S_{K}$ and parameters $t>L>1$ (with $t$ sufficiently large) such that
$\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2}.$
By Definition 5.5 and Definition 5.6, every $\mathbf{x}\in R_{t,L}(S^{\ast})$ satisfies
$\mathbf{a}_{n}(\mathbf{x})\in S^{\ast}\subset S\quad\text{for all }n\geq 1,\qquad t^{n}\leq\|\mathbf{a}_{n}(\mathbf{x})\|_{\infty}<Lt^{n}.$
Since $S^{\ast}\subset\mathcal{C}_{K}$, we also have
$\min_{1\leq j\leq d}a_{n}(x_{j})\ \geq\ \frac{1}{K}\max_{1\leq j\leq d}a_{n}(x_{j})\ =\ \frac{1}{K}\,\|\mathbf{a}_{n}(\mathbf{x})\|_{\infty}\ \geq\ \frac{t^{n}}{K}\ \longrightarrow\ \infty.$
In particular, $\mathbf{x}\in\mathcal{E}_{S}^{\mathrm{vec}}$, and hence
$R_{t,L}(S^{\ast})\subset\mathcal{E}_{S}^{\mathrm{vec}}\subset\mathcal{E}_{S}.$
Therefore
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})\ \geq\ \dim_{H}R_{t,L}(S^{\ast})\ =\ \frac{d}{2},\qquad\text{and hence}\qquad\dim_{H}(\mathcal{E}_{S})\geq\frac{d}{2}.$
Combining the upper and lower bounds, we conclude that
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=\dim_{H}(\mathcal{E}_{S})=\frac{d}{2}.$
6.2 Proof of Theorem 1.4
In the relative setting we assume throughout that the ambient set $S\subset\mathbb{N}^{d}$ has polynomial density exponent $\alpha\geq 1$ and is uniformly KK–balanced, i.e. $S\subset\mathcal{C}_{K}$ for some $K\geq 1$. Let $A\subset S$ satisfy $\overline{d}(A\mid S)>0$. Proposition 5.7 then produces an extreme seed root set $S^{\ast}\subset S$ with $\overline{d}(S^{\ast}\mid S)=0$ and parameters $t>L>1$ such that
$\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2\alpha}.$ (40)
To retain the density of $A$ while keeping the seed geometry, we will insert digits from $A\setminus S^{\ast}$ in sparse windows determined by a fast increasing sequence $(M_{k})$.
6.2.1 A density bookkeeping lemma.
The next lemma guarantees that one can choose the insertion windows so that the union of inserted digits recovers the relative density.
Lemma 6.1 (Density recovery).
Let $A_{1}\subset\mathbb{N}^{d}$ have polynomial density of exponent $\alpha\geq 1$, and let $A_{2}\subset A_{1}$ satisfy $\overline{d}(A_{2}\mid A_{1})>0$. Fix $t>L>1$. Then there exists a strictly increasing sequence $(M_{k})_{k\geq 1}$ such that, if for each $k\geq 1$ we define
$I_{k}$ $:=\bigl\{v\in\mathbb{N}^{d}:\ Lt^{M_{k-1}}+1\leq\|v\|_{\infty}\leq Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr\},$ (41)
$W_{k}$ $:=A_{2}\cap I_{k},$
then
$\overline{d}\!\left(\bigcup_{k\geq 1}W_{k}\,\Bigm|\,A_{1}\right)\;=\;\overline{d}(A_{2}\mid A_{1}).$ (42)
Proof.
Set $\delta:=\overline{d}(A_{2}\mid A_{1})>0$. By definition of $\overline{d}(\cdot\mid\cdot)$, there exists an increasing sequence $(L_{j})_{j\geq 1}$ such that
$\frac{\#(A_{2}\cap Q_{L_{j}})}{\#(A_{1}\cap Q_{L_{j}})}\longrightarrow\delta\qquad(j\to\infty).$ (43)
We inductively choose a subsequence $(L_{j_{k}})$ and then define $(M_{k})$. Set $M_{0}:=0$ and assume $M_{0},\dots,M_{k-1}$ have been fixed. Let
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$ (44)
Since $(A_{1}\cap Q_{n})\to\infty$ and $r_{k,\min}$ is fixed at stage $k$, we may choose $j_{k}$ so large that $L_{j_{k}}>r_{k,\min}+1$ and
$\left|\frac{\#(A_{2}\cap Q_{L_{j_{k}}})}{\#(A_{1}\cap Q_{L_{j_{k}}})}-\delta\right|<\frac{1}{k},\qquad\frac{\#(A_{1}\cap Q_{r_{k,\min}})}{\#(A_{1}\cap Q_{L_{j_{k}}})}<\frac{1}{k}.$ (45)
Define
$m_{k}:=L_{j_{k}}-r_{k,\min}\in\mathbb{N},\qquad M_{k}:=m_{k}^{2d}.$ (46)
Because $\|v\|_{\infty}\in\mathbb{N}$, the window in (41) coincides with the integer shell
$I_{k}=\bigl\{v\in\mathbb{N}^{d}:\ r_{k,\min}+1\leq\|v\|_{\infty}\leq r_{k,\min}+m_{k}\bigr\}=\bigl\{v\in\mathbb{N}^{d}:\ r_{k,\min}+1\leq\|v\|_{\infty}\leq L_{j_{k}}\bigr\}.$ (47)
For the lower bound in (42), evaluate the $\limsup$ at radii $R_{k}:=L_{j_{k}}$. Since $W_{k}\subset(\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}}$,
$\frac{\#((\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}.$ (48)
Moreover, by (47),
$\#(W_{k})=\#(A_{2}\cap Q_{R_{k}})-\#(A_{2}\cap Q_{r_{k,\min}})\geq\#(A_{2}\cap Q_{R_{k}})-\#(A_{1}\cap Q_{r_{k,\min}}),$ (49)
hence
$\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(A_{2}\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}-\frac{\#(A_{1}\cap Q_{r_{k,\min}})}{\#(A_{1}\cap Q_{R_{k}})}>\delta-\frac{1}{k}-\frac{1}{k}$ (50)
by (45). Taking $\limsup_{k\to\infty}$ gives $\overline{d}(\bigcup_{m\geq 1}W_{m}\mid A_{1})\geq\delta$.
For the matching upper bound, note $\bigcup_{m\geq 1}W_{m}\subset A_{2}$, so $\overline{d}(\bigcup_{m\geq 1}W_{m}\mid A_{1})\leq\overline{d}(A_{2}\mid A_{1})=\delta$. Thus (42) holds. ∎
6.2.2 Windows and a uniform product estimate.
Fix an increasing sequence $(M_{k})$ (to be chosen) and define the insertion windows
$I_{k}:=\bigl\{v\in\mathbb{N}^{d}:\ Lt^{M_{k-1}}+1\leq\|v\|_{\infty}\leq Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr\}.$ (51)
The next lemma is the bookkeeping bound needed in the almost–Lipschitz elimination step.
Lemma 6.2.
Assume $0<M_{k-1}<M_{k}$ and let $I_{k}$ be given by (51). If (32) and (34) hold, then
$\prod_{w\in I_{k}}\|w\|_{\infty}^{-2}\;\geq\;\left(\prod_{i=M_{k-1}+1}^{M_{k}}t^{-2i}\right)^{3\varepsilon_{k}}.$ (52)
Proof.
By (51) we have
$\#I_{k}=\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr)^{d}-\bigl(Lt^{M_{k-1}}\bigr)^{d}\leq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{d}.$ (53)
Condition (32) is exactly the growth hypothesis ensuring $Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\leq 2M_{k}^{1/(2d)}$ for all large $k$, hence
$\#I_{k}\leq(2M_{k}^{1/(2d)})^{d}=2^{d}M_{k}^{1/2}.$ (54)
Also, for every $w\in I_{k}$ we have $\|w\|_{\infty}\leq Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1$, so
$\prod_{w\in I_{k}}\|w\|_{\infty}^{-2}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2\#I_{k}}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2^{d+1}M_{k}^{1/2}}$ (55)
by (54). Finally, (34) is precisely the hypothesis that the last expression is bounded below by the right-hand side of (52). ∎
6.2.3 Proof of Theorem 1.4.
Proof of Theorem 1.4.
Assume $S\subset\mathbb{N}^{d}$ has polynomial density exponent $\alpha\geq 1$ and is uniformly KK–balanced, i.e. $S\subset\mathcal{C}_{K}$ for some $K\geq 1$. Let $A\subset S$ satisfy $\overline{d}(A\mid S)>0$. By Proposition 5.7, there exist a subset $S^{\ast}\subset S$ with $\overline{d}(S^{\ast}\mid S)=0$ and parameters $t>L>1$ (with $t$ sufficiently large) such that
$\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2\alpha}.$ (56)
Since $\overline{d}(S^{\ast}\mid S)=0$ and $A\subset S$, we also have $\overline{d}(A\cap S^{\ast}\mid S)=0$, hence
$\overline{d}(A\setminus S^{\ast}\mid S)=\overline{d}(A\mid S).$ (57)
Now fix an increasing sequence $(M_{k})$ and define $I_{k}$ by (51). Set
$W_{k}:=(A\setminus S^{\ast})\cap I_{k}.$ (58)
We choose $(M_{k})$ inductively so that:
  • the surgery hypotheses (32)–(35) hold (so that the almost–Lipschitz elimination applies), and
  • simultaneously, $(M_{k})$ satisfies Lemma 6.1 with $A_{1}=S$ and $A_{2}=A\setminus S^{\ast}$ (so that the inserted digits recover $\overline{d}(A\mid S)$).
At stage $k$, this is achieved by taking $M_{k}$ sufficiently large; in particular, in the proof of Lemma 6.1 one may increase the chosen radius $L_{j_{k}}$ further to meet the finitely many growth requirements at that step.
With these choices, Lemma 6.2 and Lemma 5.11 ensure (36), hence all hypotheses of Theorem 5.12 are satisfied. The elimination map is Hölder with any exponent $\gamma\in(0,1)$, so
$\dim_{H}R_{t,L}(S^{\ast},A\setminus S^{\ast},(M_{k}),(W_{k}))\geq\dim_{H}R_{t,L}(S^{\ast})=\frac{d}{2\alpha}.$ (59)
On the other hand, every point in $R_{t,L}(S^{\ast},A\setminus S^{\ast},(M_{k}),(W_{k}))$ uses only digits from $S$, because $S^{\ast}\subset S$ and $A\setminus S^{\ast}\subset S$. Therefore Proposition 5.12 yields
$\dim_{H}R_{t,L}(S^{\ast},A\setminus S^{\ast},(M_{k}),(W_{k}))\leq\dim_{H}\{x\in\mathcal{E}:\ D(x)\subset S\}=\frac{d}{2\alpha},$ (60)
where the final equality uses $S\subset\mathcal{C}_{K}$ and the sharp upper bound in the balanced case.
Combining (59) and (60) gives equality throughout, so the constructed set has Hausdorff dimension $d/(2\alpha)$.
Finally, by Lemma 6.1 applied with $A_{1}=S$ and $A_{2}=A\setminus S^{\ast}$, together with (57), the union $\bigcup_{k\geq 1}W_{k}$ satisfies
$\overline{d}\!\left(\bigcup_{k\geq 1}W_{k}\,\Bigm|\,S\right)=\overline{d}(A\setminus S^{\ast}\mid S)=\overline{d}(A\mid S).$ (61)
This is exactly the required density recovery in the relative construction, and completes the proof. ∎
6.3 Proof of Theorem 1.1
To ensure divergence, we require a lemma similar to Lemma 6.1. Since Theorem 1.1 does not include the balanced ambient condition, divergence is not automatic. However, with positive upper density, we can overcome the balanced ambient condition.
Remark 6.3.
Compared with [24], the insertion–elimination scheme is the same. The extra input in dimension $d$ is geometric and combinatorial:
  • We use the product structure and one–dimensional geometry of cylinders (Lemma 2.4 and Lemma 2.1) to control side lengths and diameters in $(0,1)^{d}$.
  • To ensure disjointness and quantitative gaps between sibling cylinders, we impose 22–separation in $\|\cdot\|_{\infty}$ and apply the separation lemmas (Lemma 2.7 and Lemma 2.6).
  • Divergence is not automatic from polynomial growth in $\mathbb{N}^{d}$. In the Banach/upper density setting we can avoid a global balance assumption by using diverging translations: Lemma 2.11 produces witnessing boxes $Q_{N_{k}}(\mathbf{v}_{k})$ with $\min_{i}(v_{k})_{i}\to\infty$, so the digits chosen from these boxes drift to infinity in every coordinate, which enforces divergence when needed.
  • Under positive upper density, bounded anisotropy can be recovered on a positive–density subset: Lemma 2.13 and Lemma 2.12 allow us to restrict to $A\cap\mathcal{C}_{K}$ with $\overline{d}(A\cap\mathcal{C}_{K})>0$, so the later covering estimates may be carried out in uniformly balanced regimes even if the original set is not balanced.
We will need a density–recovery lemma that simultaneously enforces divergence: it selects insertion blocks $W_{k}\subset A$ whose union has the same upper density as $A$, while $\min_{j}w_{j}\to\infty$ along the blocks. This will be applied in the proof of Theorem 1.1, Part (a).
Lemma 6.4 (Density recovery with divergence).
Let $A\subset\mathbb{N}^{d}$ satisfy $\overline{d}(A)=:\delta>0$. Fix $t>L>1$. Then there exists a strictly increasing sequence $(M_{k})_{k\geq 1}$ such that, if for each $k\geq 1$ we define
$I_{k}$ $:=\bigl\{v\in\mathbb{N}^{d}:\ Lt^{M_{k-1}}+1\leq\|v\|_{\infty}\leq Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr\},$ (62)
$T_{k}$ $:=\bigl\{v\in\mathbb{N}^{d}:\ \min_{1\leq j\leq d}v_{j}\geq k\bigr\},$
$W_{k}$ $:=A\cap I_{k}\cap T_{k},$
then
$\overline{d}\!\left(\bigcup_{k\geq 1}W_{k}\right)=\overline{d}(A)=\delta,$ (63)
and moreover every $w\in W_{k}$ satisfies $\min_{1\leq j\leq d}w_{j}\geq k$, hence
$\min_{1\leq j\leq d}w_{j}\xrightarrow[k\to\infty]{}\infty\qquad\text{uniformly for }w\in W_{k}.$ (64)
Proof.
For $m\geq 1$ set $A^{(m)}:=A\cap T_{m}$. Since $\mathbb{N}^{d}\setminus T_{m}=\bigcup_{i=1}^{d}\{v\in\mathbb{N}^{d}:\ v_{i}\leq m-1\}$ and
$\bigl|\{v\in Q_{N}:\ v_{i}\leq m-1\}\bigr|\leq(m-1)N^{d-1},$
we have $|(\mathbb{N}^{d}\setminus T_{m})\cap Q_{N}|/|Q_{N}|\leq d(m-1)/N\to 0$ as $N\to\infty$. Hence $\overline{d}(\mathbb{N}^{d}\setminus T_{m})=0$ and therefore
$\overline{d}(A^{(m)})=\overline{d}(A\cap T_{m})=\overline{d}(A)=\delta\qquad\text{for every fixed }m.$ (65)
We now inductively choose $(M_{k})$ together with auxiliary radii $(R_{k})$. Set $M_{0}:=0$. Assume $M_{0},\dots,M_{k-1}$ have been fixed and define
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$ (66)
Using (65) with $m=k$, we can choose $R_{k}\in\mathbb{N}$ so large that $R_{k}>r_{k,\min}+1$ and
$\left|\frac{|A^{(k)}\cap Q_{R_{k}}|}{|Q_{R_{k}}|}-\delta\right|<\frac{1}{k},\qquad\frac{|Q_{r_{k,\min}}|}{|Q_{R_{k}}|}<\frac{1}{k}.$ (67)
Set
$m_{k}:=R_{k}-r_{k,\min}\in\mathbb{N},\qquad M_{k}:=m_{k}^{2d}.$ (68)
Then $M_{k}^{1/(2d)}=m_{k}$, and since $\|v\|_{\infty}\in\mathbb{N}$ we have the identity of shells
$I_{k}=\bigl\{v\in\mathbb{N}^{d}:\ r_{k,\min}+1\leq\|v\|_{\infty}\leq r_{k,\min}+m_{k}\bigr\}=\bigl\{v\in\mathbb{N}^{d}:\ r_{k,\min}+1\leq\|v\|_{\infty}\leq R_{k}\bigr\}.$ (69)
Define $W_{k}:=A\cap I_{k}\cap T_{k}=A^{(k)}\cap I_{k}$ as in (62).
For the lower bound on $\overline{d}(\bigcup_{m\geq 1}W_{m})$, evaluate at $N=R_{k}$. Since $W_{k}\subset(\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}}$, we get
$\frac{|(\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}}|}{|Q_{R_{k}}|}\geq\frac{|W_{k}|}{|Q_{R_{k}}|}.$ (70)
Moreover, by (69),
$|W_{k}|=|A^{(k)}\cap Q_{R_{k}}|-|A^{(k)}\cap Q_{r_{k,\min}}|\geq|A^{(k)}\cap Q_{R_{k}}|-|Q_{r_{k,\min}}|.$ (71)
Dividing by $|Q_{R_{k}}|$ and using (67) yields
$\frac{|W_{k}|}{|Q_{R_{k}}|}\geq\frac{|A^{(k)}\cap Q_{R_{k}}|}{|Q_{R_{k}}|}-\frac{|Q_{r_{k,\min}}|}{|Q_{R_{k}}|}>\delta-\frac{1}{k}-\frac{1}{k}.$ (72)
Taking $\limsup_{k\to\infty}$ gives $\overline{d}(\bigcup_{m\geq 1}W_{m})\geq\delta$.
For the reverse inequality, note $\bigcup_{m\geq 1}W_{m}\subset A$, hence $\overline{d}(\bigcup_{m\geq 1}W_{m})\leq\overline{d}(A)=\delta$. This proves (63). Finally, (64) follows immediately from $W_{k}\subset T_{k}$. ∎
6.3.1 Proof of Theorem 1.1, Part (a)
Proof of Theorem 1.1, Part (a).
Fix $S\subset\mathbb{N}^{d}$ with $\overline{d}(S)=:\delta>0$. We construct a set $\mathcal{F}_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ such that
$\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)=\delta\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S},$ (73)
and $\dim_{H}(\mathcal{F}_{S})=d/2$.
Fix $K_{0}\geq 2$ and set $S_{0}:=\mathcal{C}_{K_{0}}$. Since $|S_{0}\cap Q_{N}|\asymp N^{d}$, the set $S_{0}$ has polynomial density exponent $1$. Applying Proposition 5.7 to $S_{0}$ (with $\alpha=1$), we obtain parameters $t>L>1$ (with $t$ sufficiently large) and a subset $B\subset S_{0}$ such that
$\mathcal{R}:=R_{t,L}(B)$ (74)
is non-empty and
$\dim_{H}(\mathcal{R})=\frac{d}{2}.$ (75)
Moreover, every $\mathbf{y}\in\mathcal{R}$ satisfies $\mathbf{a}_{n}(\mathbf{y})\in B\subset\mathcal{C}_{K_{0}}$ and $t^{n}\leq\|\mathbf{a}_{n}(\mathbf{y})\|_{\infty}<Lt^{n}$, hence
$\min_{1\leq j\leq d}a_{n}(y_{j})\geq\frac{1}{K_{0}}\,\|\mathbf{a}_{n}(\mathbf{y})\|_{\infty}\geq\frac{t^{n}}{K_{0}}\longrightarrow\infty.$ (76)
In particular, $\mathcal{R}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$.
Apply Lemma 6.4 to $A:=S$ with the above parameters $t>L>1$. We obtain a strictly increasing sequence $(M_{k})_{k\geq 1}$ and finite sets $W_{k}\subset S$ such that, writing $W:=\bigcup_{k\geq 1}W_{k}$,
$\overline{d}(W)=\overline{d}(S)=\delta,\qquad\min_{1\leq j\leq d}w_{j}\geq k\ \ \text{for all }w\in W_{k}.$ (77)
By enlarging $M_{k}$ if necessary, we may assume in addition that the growth conditions (32)–(36) required for the surgery/elimination construction hold; at each stage this imposes only finitely many constraints that are preserved when $M_{k}$ increases.
Define
$\mathcal{F}_{S}:=R_{t,L}\bigl(B,\mathbb{N}^{d},(M_{k}),(W_{k})\bigr).$ (78)
By construction, the block $W_{k}$ is inserted (in a fixed order) at the digit positions $\{M_{k},\dots,M_{k}+|W_{k}|-1\}$, independently of the choice of $\mathbf{x}\in\mathcal{F}_{S}$. Consequently,
$W\subset D(\mathbf{x})\cap S\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S}.$ (79)
Since $D(\mathbf{x})\cap S\subset S$, we obtain
$\overline{d}(S)=\overline{d}(W)\leq\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)\leq\overline{d}(S),$ (80)
and hence $\overline{d}(D(\mathbf{x})\cap S)=\overline{d}(S)$ for all $\mathbf{x}\in\mathcal{F}_{S}$.
Let $\mathbf{x}\in\mathcal{F}_{S}$. Along the background positions, the digit vectors of $\mathbf{x}$ are taken from $B\subset\mathcal{C}_{K_{0}}$ at scales comparable to $t^{n}$; therefore (76) implies coordinatewise divergence along those indices. Along the inserted positions belonging to the $k$th inserted block, every digit vector lies in $W_{k}$, hence satisfies $\min_{j}a_{n}(x_{j})\geq k$ by (77). Since $k\to\infty$ as we move to later blocks, the inserted digits also diverge coordinatewise. Hence $\mathbf{a}_{n}(\mathbf{x})\to\infty$ coordinatewise, and therefore
$\mathcal{F}_{S}\subset\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}.$ (81)
Let $f:\mathcal{F}_{S}\to\mathcal{R}$ be the elimination map deleting all inserted blocks. By Theorem 5.12, $f$ is Hölder for every $\gamma\in(0,1)$ and is surjective. Hence
$\dim_{H}(\mathcal{F}_{S})\geq\dim_{H}(\mathcal{R})=\frac{d}{2}.$ (82)
Together with (81) and $\dim_{H}(\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}})=d/2$, we conclude that
$\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}:\ \overline{d}(D(\mathbf{x})\cap S)=\overline{d}(S)\Bigr\}=\frac{d}{2},$
as desired. ∎
6.3.2 Proof of Theorem 1.1, Part (b)
We will use Lemma 2.11 to convert Banach density information into deterministic insertion blocks: it replaces $S$ by $S^{\prime}\subset S$ with $\overline{d}(S^{\prime})=0$ and $d^{\ast}(S^{\prime})=d^{\ast}(S)$ and provides witnesses $Q_{N_{k}}(\mathbf{v}_{k})$ with $\min_{i}(v_{k})_{i}\to\infty$. Setting $W_{k}:=S^{\prime}\cap Q_{N_{k}}(\mathbf{v}_{k})$ gives dense blocks whose digits drift to infinity in every coordinate.
Compared with the one–dimensional argument in [24], this is the additional multidimensional step: in $\mathbb{N}^{d}$ coordinatewise divergence is not automatic (cf. $\mathbb{N}\times F$), and Lemma 2.11 enforces it by pushing the Banach witnesses off to infinity.
Proof of Theorem 1.1, Part (b).
Let $\beta:=d^{\ast}(S)>0$. If $d^{\ast}(S)=\overline{d}(S)$, then Part (a) applied to $S$ yields a set $F_{S}$ with $\dim_{H}F_{S}=d/2$ and $\overline{d}(D(\mathbf{x})\cap S)=\overline{d}(S)=\beta$ for every $\mathbf{x}\in F_{S}$. Since $d^{\ast}(\cdot)\geq\overline{d}(\cdot)$ and $D(\mathbf{x})\cap S\subset S$, we also have $d^{\ast}(D(\mathbf{x})\cap S)=\beta=d^{\ast}(S)$, proving Part (b).
Assume henceforth that $d^{\ast}(S)>\overline{d}(S)$. Apply Lemma 2.11 to obtain a subset $S_{0}\subset S$, a strictly increasing sequence $(N_{k})_{k\geq 1}$, and vectors $(\mathbf{v}_{k})_{k\geq 1}\subset\mathbb{N}^{d}$ such that $d^{\ast}(S_{0})=d^{\ast}(S)=\beta$, $\overline{d}(S_{0})=0$, $\min_{i}(v_{k})_{i}\to\infty$, and
$\frac{|S_{0}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}\longrightarrow\beta.$ (83)
Replacing $S$ by $S_{0}$ (and relabelling), we may assume $\overline{d}(S)=0$ and (83) holds with $\beta=d^{\ast}(S)$.
Passing to a subsequence, we may assume the boxes $B_{k}:=Q_{N_{k}}(\mathbf{v}_{k})$ are pairwise disjoint. Define $W_{k}:=S\cap B_{k}$ and $m_{k}:=|W_{k}|$. Then (83) becomes
$\frac{|W_{k}|}{|Q_{N_{k}}|}\longrightarrow d^{\ast}(S)=\beta.$ (84)
Moreover, since $W_{k}\subset\mathbf{v}_{k}+Q_{N_{k}}$, every $w\in W_{k}$ satisfies $\min_{i}w_{i}\geq\min_{i}(v_{k})_{i}+1$, hence $\min_{i}w_{i}\to\infty$ uniformly for $w\in W_{k}$.
Since $\overline{d}(S)=0$, the complement $S^{c}:=\mathbb{N}^{d}\setminus S$ has upper density $1$. By Lemma 2.13, there exists $K\geq 1$ such that $\overline{d}(S^{c}\cap\mathcal{C}_{K})>0$, where $\mathcal{C}_{K}:=\{\mathbf{a}\in\mathbb{N}^{d}:\max_{j}a_{j}/\min_{j}a_{j}\leq K\}$. Set $A:=S^{c}\cap\mathcal{C}_{K}$. Then $A$ is uniformly KK–balanced and has polynomial density exponent $1$.
Applying Proposition 5.7 to $A$, we obtain parameters $t>L>1$ (with $t$ sufficiently large) and an infinite 22–separated set $B\subset A$ such that $\mathcal{R}:=R_{t,L}(B)$ is non-empty and $\dim_{H}\mathcal{R}=d/2$. Since $B\subset\mathcal{C}_{K}$, every $\mathbf{y}\in\mathcal{R}$ satisfies $t^{n}\leq\|\mathbf{a}_{n}(\mathbf{y})\|_{\infty}<Lt^{n}$ and $\min_{j}a_{n}(y_{j})\geq\|\mathbf{a}_{n}(\mathbf{y})\|_{\infty}/K\to\infty$, so $\mathcal{R}\subset E_{\mathbb{N}^{d}}^{\mathrm{vec}}$.
Fix a decreasing sequence $(\varepsilon_{k})_{k\geq 1}$ with $\varepsilon_{k}\downarrow 0$. We choose a strictly increasing sequence $(M_{k})_{k\geq 1}$ with $M_{0}=0$ such that, for every $k\geq 1$, the data $\bigl(t,L;\,A:=S,\ B;\,(M_{i}),\ (W_{i}),\ (\varepsilon_{i})\bigr)$ satisfy (35) and (36), and in addition
$M_{k}\geq M_{k-1}+m_{k-1}+1.$ (85)
This is possible because, at stage $k$, (35)–(36) impose only finitely many explicit inequalities involving the previously chosen data and the new integer $M_{k}$, and each such inequality is preserved when $M_{k}$ is increased.
Define
$F_{S}:=R_{t,L}\bigl(B,S,(M_{k}),(W_{k})\bigr).$ (86)
By construction, every inserted digit lies in $S$, while every background digit lies in $B\subset S^{c}$, so the two digit alphabets are disjoint. Since the boxes $B_{k}$ are disjoint, the sets $W_{k}$ are disjoint as well, and $B$ is 22–separated; hence no digit vector is repeated in the construction.
We next verify divergence. Background digits come from $\mathcal{R}\subset E_{\mathbb{N}^{d}}^{\mathrm{vec}}$ and therefore diverge coordinatewise along the background positions. Inserted digits belong to $W_{k}$ and satisfy $\min_{i}w_{i}\to\infty$ along the inserted blocks. Hence every $\mathbf{x}\in F_{S}$ has coordinatewise digits diverging, and thus $F_{S}\subset E_{\mathbb{N}^{d}}^{\mathrm{vec}}$.
Let $f:F_{S}\to\mathcal{R}$ be the elimination map deleting all inserted blocks. By Theorem 5.12, $f$ is Hölder for every $\gamma\in(0,1)$ and surjective. Hence $\dim_{H}F_{S}\geq\dim_{H}\mathcal{R}=d/2$. Together with $F_{S}\subset E_{\mathbb{N}^{d}}^{\mathrm{vec}}$ and $\dim_{H}(E_{\mathbb{N}^{d}}^{\mathrm{vec}})=d/2$, we obtain
$\dim_{H}F_{S}=\frac{d}{2}.$ (87)
Finally set
$D_{*}:=\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in F_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}.$
For each $k$, the construction (86) inserts the block $W_{k}$ (in a fixed order) at the digit positions $\{M_{k},\dots,M_{k}+m_{k}-1\}$, independently of $\mathbf{x}\in F_{S}$. Thus $W_{k}\subset D_{*}\cap S$ for every $k$, and therefore
$\frac{|(D_{*}\cap S)\cap B_{k}|}{|Q_{N_{k}}|}\geq\frac{|W_{k}|}{|Q_{N_{k}}|}.$
Taking $\limsup$ and using (84) gives $d^{\ast}(D_{*}\cap S)\geq d^{\ast}(S)$, while $d^{\ast}(D_{*}\cap S)\leq d^{\ast}(S)$ is immediate from $D_{*}\cap S\subset S$. Hence
$d^{\ast}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in F_{S}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap S\Bigr)=d^{\ast}(S),$
as required. The “in particular” statement follows from (87). ∎
Acknowledgements
We are grateful to Rongzhong Xiao, Leiye Xu, and Yubin He for many helpful discussions, comments, and valuable suggestions, and for remarks that improved the exposition. This work is supported by the National Key R&D Program of China (Nos. 2024YFA1013602 and 2024YFA1013600).
Zhuowen Guo
School of Mathematical Sciences, University of Science and Technology of China
Kangbo Ouyang
School of Mathematical Sciences, University of Science and Technology of China
Jiahao Qiu
School of Mathematical Sciences, University of Science and Technology of China
Shuhao Zhang
School of Mathematical Sciences, University of Science and Technology of China